### v1

# Estimating spectral radius of Bell-type operator via finite dimensional approximation of orthogonal projections

Abstract
In quantum mechanics, the following problem is important: For two orthogonal projections $P$ and $Q$ on a separable infinite dimensional Hilbert space $V$, estimate the spectral radius $\rho([A,B])$, where $A=2P-\operatorname{id}_{V},B=2Q-\operatorname{id}_{V}$. This problem is known to be equivalent to determining the spectral radius of the Bell-CHSH operator.
To approach this problem, we give a method to approximate orthogonal projections on $V$ by those on finite dimensional subspaces.
This method provides an upper bound and a lower bound for $\rho([A,B])$, which become exact when the matrix representations of $P$ and $Q$ are “constant-angle one-shifted form”.
keywords:
Orthogonal projections , Infinite dihedral group , Free product of cyclic groups of order 2 , Bell-CHSH operator , Tsirelson’s inequality , Device-independent quantum key distribution , Tridiagonal matrix
[…]
organization=Mitsubishi Electric Corporation, Information Technology R&D Center,addressline=5-1-1 Ofuna, city=Kamakura-shi, postcode=247-8501, state=Kanagawa, country=Japan
1  Introduction
The problem of studying properties of orthogonal projections on a separable Hilbert space $V$ arises in various fields, including quantum mechanics, and has been extensively studied from a mathematical perspective [1].
For instance, given two orthogonal projections $P$ and $Q$ on $V$, let $A=2P-\operatorname{id}_{V},B=2Q-\operatorname{id}_{V}$, and consider the problem of computing the spectral radius $\rho([A,B])$ of the commutator $[A,B]$, defined by
$\rho([A,B])=\sup_{u\in V,\|u\|=1}|([A,B]u,u)|.$
This problem is known to be equivalent to determining the spectral radius of the Bell-CHSH operator [2], which is a bounded self-adjoint operator on a Hilbert tensor space $V_{1}\otimes V_{2}$ used in quantum mechanics, as shown in [3, 4]. For the definition of the Bell-CHSH operator and the background knowledge assumed in this paper, see Section 2.3.
The problem of determining the spectral radius of the Bell-CHSH operator has important applications, such as testing the violation of local realism in quantum mechanics [5], and evaluating the security of device-independent quantum key distribution protocols [6]. Owing to these applications, the computation of $\rho([A,B])$ is important.
When $V$ is finite dimensional, this problem can be solved immediately by the existing results. First, it is known that there is a one-to-one correspondence between pairs of orthogonal projections on $V$ and unitary representations of the infinite dihedral group $D_{\infty}$ on $V$. Using group representation theory, Behncke [7] showed that, regardless of whether $V$ is finite or infinite dimensional, $V$ is irreducible only if $\dim V\leq 2$. Here, $V$ is said to be irreducible if every invariant subspace under $P$ and $Q$ is trivial.
By combining this result with the general theory of decomposition of group representations into irreducible components, it follows that if $V$ is finite dimensional, then $V$ can be decomposed as an orthogonal direct sum of invariant subspaces under $P$ and $Q$, each of which is at most two dimensional. In quantum information theory, this fact is known as Jordan’s lemma [8, 9].
Using this decomposition, the problem of computing $\rho([A,B])$ in finite dimensional $V$ reduces to the case where $\dim V\leq 2$. In these cases, it is known that $P$ and $Q$ admit a simple matrix representation using Halmos’ theorem [10]. Based on this representation and elementary calculations, we find that $\rho([A,B])$ can be computed using the eigenvalues of $A+B$.
From these observations, we conclude that if $V$ is finite dimensional, let $\lambda$ be the eigenvalue of $A+B$ whose absolute value is closest to $\sqrt{2}$, then the following holds:
$\rho([A,B])=b(\lambda),$ (1.1)
where $b$ is defined by $b(\lambda)=2\sqrt{1-\left(\dfrac{\lambda^{2}}{2}-1\right)^{2}}\quad(\lambda\in\mathbb{R})$.
On the other hand, when $V$ is infinite dimensional, it is not known whether or how equation (1.1) can be extended, either in the form of an equality or an inequality. One reason for this is the existence of pairs of orthogonal projections on an infinite dimensional Hilbert space that admit no nonzero finite dimensional invariant subspaces under both projections [11].
In this paper, we consider the problem of extending equation (1.1) in the form of an inequality as a representative example. As in this case, there are several known problems concerning the properties of orthogonal projections that have only been resolved in the finite dimensional case.
This work makes three main contributions. First, we prove that every finite family of orthogonal projections on an infinite dimensional Hilbert space can be approximated, in the sense of Theorem 1.1, by orthogonal projections on finite dimensional subspaces. For details of this result, see Section 3.1. This approximation allows certain inequalities to be reduced to the finite dimensional case. This method is stated in Section 3.2.
As a simple application, in Section 3.3, we present a new proof of Tsirelson’s inequality [12], which asserts that the spectral radius of every Bell-CHSH operator is bounded above by $2\sqrt{2}$. We compare our proof with others. When either $V_{1}$ or $V_{2}$ has dimension at most two, Tsirelson’s inequality can be proved using elementary calculus. Tsirelson [13] showed that this inequality can be reduced to the case of dimension at most two by an ingenious technique using the Clifford algebra. Based on our approximation method and Jordan’s lemma, we reduce the proof to this case.
Second, for two orthogonal projections $P$ and $Q$, we provide an explicit formula for a finite dimensional approximation of the pair $(P,Q)$. This can be summarized as follows. We prove the existence of an orthogonal direct sum decomposition of $V$ into invariant subspaces under $P$ and $Q$, in which every infinite dimensional subspace admits an explicit finite dimensional approximation of the pair $(P,Q)$ as described in Theorem 1.1. We refer to the infinite dimensional matrix representations admitting this approximation as the one-shifted form. For the definition of one-shifted form, see Definition 4.1. From the viewpoint of group representation theory, two orthogonal projections having a one-shifted form yield a cyclic unitary representation of $D_{\infty}$. Theorem 1.1 thus provides an explicit formula that decomposes every infinite dimensional unitary representation of $D_{\infty}$ into cyclic representations. In other words, the one-shifted form (Definition 4.1) extends Jordan’s lemma to the infinite dimensional case. For details of this theorem, see Section 4.
Theorem 1.1 (Approximation of one-shifted form).
There exists an orthonormal basis $\mathcal{U}=\{u_{i}\}_{i=1}^{\infty}$ and a pair of sequences $\Theta=\{\theta_{i}\}_{i=1}^{\infty},\Omega=\{\omega_{i}\}_{i=1}^{\infty}\subset(0,\pi)$ such that, for any vector $v\in V$,
$\max\{\|v-\pi_{n}v\|,\|Pv-P_{n}(\Theta)\pi_{n}v\|,\|Qv-Q_{n}(\Omega)\pi_{n}v\|\}\to 0\quad(n\to\infty),$
where $\pi_{n}$ is the orthogonal projection onto the finite dimensional subspace $V_{n}$ spanned by $\{u_{1},\dots,u_{2n}\}$, and $P_{n}(\Theta)$, $Q_{n}(\Omega)$ are orthogonal projections on $V_{n}$.
Moreover, the matrix representations of $P_{n}(\Theta)$ and $Q_{n}(\Omega)$ in this orthonormal basis are symmetric tridiagonal matrices given by:
[…]
where $r$ is defined by $r(x)=\begin{pmatrix}\cos x&\sin x\\ \sin x&-\cos x\end{pmatrix}\quad(x\in\mathbb{R})$.
Thus we can reduce the problem of estimating $\rho([A,B])$ to the case where $P$ and $Q$ admit a one-shifted form.
Third, when $P$ and $Q$ admit a one-shifted form, we derive the inequalities as stated in Theorem 1.2. This theorem estimates the spectral radius of $[A,B]$ from above and below using the eigenvalues of symmetric tridiagonal matrices approximating $A+B$, based on the explicit formula given in Theorem 1.1. Theorem 1.2 is an extension of equation (1.1) to infinite dimensional Hilbert spaces.
Theorem 1.2.
Assume that $P$ and $Q$ admit the approximation as in Theorem 1.1. Then we have
$\sup_{\lambda\in F(A+B)\cup\{0\}}b(\lambda)\leq\rho([A,B])\leq\liminf_{n\to\infty}b(\lambda_{n}),$ (1.2)
where for $n\in\mathbb{N}$, we define $A_{n}=2P_{n}-\operatorname{id}_{V_{n}},B_{n}=2Q_{n}-\operatorname{id}_{V_{n}}$, and let $\lambda_{n}$ be the eigenvalue of $A_{n}+B_{n}$ whose absolute value is closest to $\sqrt{2}$.
The set $F(A+B)$ consists of those $\lambda\in(0,2)$ for which there exist a sequence $\{\eta_{k}\}_{k=1}^{\infty}$ and vectors $v_{-,k},v_{+,k}\in V_{k}$ $(k=1,2,\dots)$ satisfying the following conditions:
  1. (i)
    For any $k\in\mathbb{N}$, $\eta_{k}$ and $\eta_{k}$ are eigenvalues of the symmetric tridiagonal matrix $A_{k}+B_{k}$, and $v_{-,k},v_{+,k}$ are unit eigenvectors corresponding to $\eta_{k}$ and $\eta_{k}$, respectively.
  2. (ii)
    $\lim_{k\to\infty}\eta_{k}=\lambda,\quad\lim_{k\to\infty}\max\{\|(P-P_{k})v_{\pm,k}\|,\|(Q-Q_{k})v_{\pm,k}\|\}=0$.
These inequalities are useful because there exist concrete examples in which equality holds. One such case, in which both $\Theta$ and $\Omega$ are equal to a constant $\theta$ in the one-shifted form, is referred to as a constant-angle one-shifted form.
In this setting, we can compute the exact value of $\rho([A,B])$ using Theorem 1.2 and known results for symmetric tridiagonal matrices:
$\rho([A,B])=\left\{\begin{array}[]{ll}2|\sin(2\theta)|&\quad\left|\theta-\dfrac{\pi}{2}\right|>\dfrac{\pi}{4},\vskip 5.69054pt\\ 2&\quad\left|\theta-\dfrac{\pi}{2}\right|\leq\dfrac{\pi}{4}.\end{array}\right$.
For the results on symmetric tridiagonal matrices used in this paper, see Section 2.4.
As shown in Section 4.2, there exists an example in which the orthogonal projections $P$ and $Q$ have such a one-shifted form but do not admit nonzero finite dimensional invariant subspaces under $P$ and $Q$. For instance, this occurs if $\theta=\pi/2$.
2  Preliminaries
2.1 Notations
Throughout this paper, we adopt the following notation.
  • 1.
    $\mathbb{Z}_{2}:\mathbb{Z}/2\mathbb{Z}$.
  • 2.
    $\Gamma_{m}:$ The free product of $m$ copies of $\mathbb{Z}_{2}$, often denoted as $\underbrace{\mathbb{Z}_{2}*\cdots*\mathbb{Z}_{2}}_{\text{$m$ times}}$ for $m\in\mathbb{N}$. Elements $g_{i}$ are the generators of the $i$-th copy of $\mathbb{Z}_{2}$.
  • 3.
    $D_{\infty}:$ $\Gamma_{2}$, also known as the infinite dihedral group.
  • 4.
    $\tau=\tau_{P_{1},\dots,P_{m}}:$ The unitary representation defined by Fact 2.1, for orthogonal projections $P_{1},\dots,P_{m}$ on a Hilbert space $V$.
  • 5.
    $\operatorname{id}_{V}:$ The identity map on a Hilbert space $V$.
  • 6.
    $E_{n}:$ The identity matrix on $\mathbb{C}^{n}$.
  • 7.
    $\mathbb{C}\langle x_{1},\dots,x_{m}\rangle:$ The set of all non-commutative polynomials over $\mathbb{C}$ with non-commutative variables $x_{1},\dots,x_{m}$.
  • 8.
    $\bigoplus_{i\in I}T_{i}:$ The operator on a direct sum $\bigoplus_{i\in I}V_{i}$ of Hilbert spaces $\{V_{i}\}_{i\in I}$ for a family $\{T_{i}\}_{i\in I}$ of bounded linear operators where each $T_{i}$ acts on $V_{i}$, the action is defined by
    $\bigoplus_{i\in I}T_{i}\sum_{i\in I}v_{i}=\sum_{i\in I}T_{i}v_{i}\quad(i\in I,v_{i}\in V_{i}).$
  • 9.
    $\sigma(T):$ The spectrum for a bounded linear operator $T$ on a Hilbert space $V$.
  • 10.
    $\sigma_{p}(T):$ The set of all eigenvalues of a bounded linear operator $T$ on a Hilbert space $V$.
  • 11.
    $\rho(T):$ The spectral radius of a bounded normal operator $T$ on a Hilbert space $V$. $\rho(T)$ is defined by
    $\rho(T)=\sup_{v\in V,\|v\|=1}|(Tv,v)|.$
  • 12.
    $\bigotimes_{i=1}^{m}V_{i}=V_{1}\otimes\cdots\otimes V_{m}:$ The Hilbert tensor product of separable Hilbert spaces $V_{1},\cdots,V_{m}$.
  • 13.
    $P_{i}:$ The abbreviation for $\operatorname{id}_{V_{1}}\otimes\dots\otimes\operatorname{id}_{V_{i-1}}\otimes P_{i}\otimes\operatorname{id}_{V_{i+1}}\otimes\dots\otimes\operatorname{id}_{V_{m}}$ acting on a Hilbert tensor product $\bigotimes_{i=1}^{m}V_{i}$, where $P_{i}$ is an orthogonal projection on $V_{i}$.
2.2 Unitary representations of $\mathbb{Z}_{2}*\cdots*\mathbb{Z}_{2}$
For a given Hilbert space $V$ and $m\in\mathbb{N}$, there exists a one-to-one correspondence between $m$ orthogonal projections on $V$ and unitary representations of $\Gamma_{m}$ on $V$, as described below. For basic notions and properties of unitary representations of groups, see, e.g., [14].
Fact 2.1.
Let $P_{1},\dots,P_{m}$ be $m$ orthogonal projections on $V$, and define the representation $\tau=\tau_{P_{1},\dots,P_{m}}$ of $\Gamma_{m}$ by $\tau(g_{i})=2P_{i}-\operatorname{id}_{V}\quad(i=1,\dots,m)$. Then $(V,\tau)$ is a unitary representation of $\Gamma_{m}$.
Conversely, for a unitary representation $\tau$ of $\Gamma_{m}$ whose representation space is $V$, define $P_{1},\dots,P_{m}$ by $P_{i}=\dfrac{\tau(g_{i})+\operatorname{id}_{V}}{2}\quad(i=1,\dots,m)$. Then $P_{1},\dots,P_{m}$ are orthogonal projections on $V$.
The following Fact 2.2 is well known concerning the irreducible decomposition of finite dimensional unitary representations of $D_{\infty}$. In quantum information theory, Fact 2.2 is referred to as Jordan’s lemma.
Fact 2.2 (Jordan’s lemma).
Let
  1. (S1)
    $V$ be a separable Hilbert space,
  2. (S2)
    $P$ and $Q$ be a pair of orthogonal projections on $V$, and
  3. (S3)
    AA and $B$ be defined by $A=2P-\operatorname{id}_{V}$ and $B=2Q-\operatorname{id}_{V}$,
then
  1. (i)
    If $\dim V<\infty$, there exists $M\in\mathbb{N}$ and an orthogonal direct sum decomposition $V=\bigoplus_{i=1}^{M}W_{i}$ such that $W_{i}$ is an irreducible invariant subspace under $\tau_{P,Q}$ for any $i\in\{1,\dots,M\}$.
  2. (ii)
    If $\tau_{P,Q}$ is an irreducible unitary representation of $D_{\infty}$, then $\dim V\leq 2$.
  3. (iii)
    If $\tau_{P,Q}$ is an irreducible unitary representation of $D_{\infty}$ and $\dim V=2$, then there exists $x\in(0,1)$ such that with respect to a suitable orthonormal basis
    $2P-E_{2}=\begin{pmatrix}2x-1&2\sqrt{x(1-x)}\\ 2\sqrt{x(1-x)}&-(2x-1)\end{pmatrix},2Q-E_{2}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$.
Proof of (i).
This follows from the general theory of finite dimensional unitary representations of groups; see, e.g., [14]. ∎
Proof of (ii).
This was proven by Behncke [7]. ∎
Proof of (iii).
This follows from Halmos’ theorem [10]. ∎
The following result is a consequence of Fact 2.2 and elementary calculations.
Fact 2.3.
Let
  1. (S1)
    $V$ be a finite dimensional Hilbert space,
  2. (S2)
    $P$ and $Q$ be a pair of orthogonal projections on $V$, and
  3. (S3)
    $A,B$ be defined by $A=2P-\operatorname{id}_{V}$ and $B=2Q-\operatorname{id}_{V}$,
then
  1. (i)
    If $\dim V=1$, then we have
    $\sigma(A+B)\subset\{-2,0,2\},\sigma([A,B])=\{0\}.$
  2. (ii)
    If $\dim V=2$ and $\tau_{P,Q}$ is an irreducible unitary representation of $D_{\infty}$, then there exists $x\in(0,1)$ such that
    $\sigma(A+B)$ $=\{-2\sqrt{x},2\sqrt{x}\},\sigma(A-B)=\{-2\sqrt{1-x},2\sqrt{1-x}\},$
    $\sigma([A,B])$ $=\{-4\sqrt{x(1-x)}i,4\sqrt{x(1-x)}i\}.$
    Define $u(x),v(x)$ by
    $u(x)$ $=\dfrac{1}{\sqrt{\dfrac{1-x}{(1+\sqrt{x})^{2}}+1}}\begin{pmatrix}-\dfrac{\sqrt{1-x}}{1+\sqrt{x}}\\ 1\end{pmatrix}$,
    $v(x)$ $=\dfrac{1}{\sqrt{\dfrac{1-x}{(1-\sqrt{x})^{2}}+1}}\begin{pmatrix}\dfrac{\sqrt{1-x}}{1-\sqrt{x}}\\ 1\end{pmatrix}$.
    Then $u(x),v(x)$ are unit eigenvectors of $A+B$, corresponding to
    $2\sqrt{x},2\sqrt{x}$, respectively.
  3. (iii)
    For any orthogonal projections $P,Q$ in a two dimensional Hilbert space $V$ such that $\tau_{P,Q}$ is an irreducible unitary representation of $D_{\infty}$,
    $\rho([A,B])=|([A,B](\alpha u(x)+\beta v(x)),\alpha u(x)+\beta v(x))|,$
    for some $\alpha,\beta\in\mathbb{C}\text{ such that }|\alpha|^{2}+|\beta|^{2}=1$.
  4. (iv)
    Let $\lambda$ be the eigenvalue of $A+B$ whose absolute value is closest to $\sqrt{2}$. Then
    $\sigma([A,B])$ $=b(\lambda).$
2.3 Bell-CHSH operator
2.3.1 Spectral radius of Bell-CHSH operator
We define the Bell-CHSH operator and briefly summarize various results on it. For two separable Hilbert spaces $V_{1}$ and $V_{2}$, and a pair of orthogonal projections $P_{i,1}$ and $P_{i,2}$ on $V_{i}$ for $i=1,2$, define $A_{j}=2P_{1,j}-\operatorname{id}_{V_{1}}$ and $B_{j}=2P_{2,j}-\operatorname{id}_{V_{2}}$ for $j=1,2$. The Bell-CHSH operator $\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})$ on $V_{1}\otimes V_{2}$ is defined as follows:
$\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})=(A_{1}+A_{2})\otimes B_{1}+(A_{1}-A_{2})\otimes B_{2}.$
We denote $\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})$ by $\mathcal{B}$ whenever no confusion arises.
The spectral radius of $\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})$ is given by the following formula as established in [3, 4],
$\rho(\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2}))=\sqrt{4+\rho([A_{1},A_{2}])\rho([B_{1},B_{2}])}.$ (2.1)
This formula shows that the problem of estimating $\rho(\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2}))$ is equivalent to that of estimating $\rho([A_{1},A_{2}])$ and $\rho([B_{1},B_{2}])$. Remark that
$A_{1},A_{2}]=[P_{1,1},P_{1,2}],[B_{1},B_{2}]=[P_{2,1},P_{2,2}].$
2.3.2 Tsirelson’s inequality
Tsirelson’s inequality is stated as Theorem 2.4.
Theorem 2.4 (Tsirelson’s inequality [12]).
Let
  1. (S1)
    $V_{1}$ and $V_{2}$ be separable Hilbert spaces,
  2. (S2)
    $P_{i,1}$ and $P_{i,2}$ be orthogonal projections on $V_{i}$ for $i\in\{1,2\}$,
then
$\rho\left(\mathcal{B}\left(P_{1,1},P_{1,2},P_{2,1},P_{2,2}\right)\right)\leq 2\sqrt{2}.$ (2.2)
2.4 Tridiagonal matrices
2.4.1 Sufficient conditions for distinct eigenvalues
We apply the following well-known fact to estimate the spectral radius of $[A,B]$; see, e.g., [15] for the details.
Fact 2.5.
Let $C$ be a symmetric tridiagonal matrix defined by
$C=\begin{pmatrix}c_{1,1}&c_{1,2}&0&0&\dots&0&0&0\\ c_{2,1}&c_{2,2}&c_{2,3}&0&\dots&0&0&0\\ 0&c_{3,2}&c_{3,3}&c_{3,4}&\dots&0&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&0&0&\dots&c_{n-1,n-2}&c_{n-1,n-1}&c_{n-1,n}\\ 0&0&0&0&\dots&0&c_{n,n-1}&c_{n,n}\end{pmatrix}$
such that
$c_{i,i+1}=c_{i+1,i}\neq 0\quad(i=1,2,\dots,n-1).$
Then the eigenvalues of $C$ are distinct.
2.4.2 Explicit eigenvalues and eigenvectors of Toeplitz matrices
Yueh [16] studied the explicit formula for eigenvalues of the following $n\times n$ tridiagonal matrix:
$X=\begin{pmatrix}b+\gamma&c&0&\dots&0&\alpha\\ a&b&c&\dots&0&0\\ 0&a&b&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&0&\dots&b&c\\ \beta&0&0&\dots&a&b+\delta\\ \end{pmatrix}$
where $n\in\mathbb{N}$ and $a,b,\gamma,\delta,\alpha,\beta\in\mathbb{C}$. Define $\mu$ by $\mu=\sqrt{\dfrac{a}{c}}$. According to Theorem 3.1 of [16], for any eigenvalue $\lambda$ of $X$ there exists $\phi\in[0,2\pi)$ such that
$\lambda=\left\{\begin{array}[]{ll}b+2c\mu\cos\phi&\quad\sin\phi\neq 0,\\ b\pm 2c\mu&\quad\sin\phi=0,\end{array}\right$.
and if $\sin\phi\neq 0$ then $\phi$ satisfies the following equation:
$\mu^{n}\left(ac\sin\left(\left(n+1\right)\phi\right)+\left(\gamma\delta-\alpha\beta\right)\sin\left(\left(n-1\right)\phi\right)-c\mu\left(\gamma+\delta\right)\sin\left(n\phi\right)\right)$
$-\left(c\alpha\mu^{2n}+\alpha\beta\right)\sin\phi=0.$
Moreover, according to (3.3) of [16], for any eigenvalue $\lambda$ of $X$ such that $\sin\phi\neq 0$, and for any eigenvector $u$ of $X$ corresponding to $\lambda$,
$u_{j}=$ $\dfrac{2i}{\sqrt{\omega}}\left\{cu_{1}\mu^{j}\sin(j\phi)-(\alpha u_{n}+\gamma u_{1})\mu^{j-1}\sin((j-1)\phi)\right\}$
$-\dfrac{2i}{\sqrt{\omega}}H_{j}^{n}(\beta u_{1}+\delta u_{n})\mu^{j-n}\sin((j-n)\phi)\quad(j=1,2,\dots,n),$ (2.3)
where
$\omega$ $=(b-\lambda)^{2}-4ac=4ac\cos^{2}\phi-4ac=-4ac\sin^{2}\phi,$
$H_{j}^{n}$ $=\left\{\begin{array}[]{ll}0&\quad j<n,\\ 1&\quad j\geq n.\end{array}\right$.
3  Finite dimensional approximation of orthogonal projections
We define finite dimensional approximations of a finite family of orthogonal projections, and present their basic properties and applications.
In Section 3.1, we define finite dimensional approximations of a finite family of orthogonal projections, and present their basic properties. In particular, we prove that for any finite family of orthogonal projections on a separable Hilbert space $V$ there exists a finite dimensional approximation.
In Section 3.2, we present an application of finite dimensional approximations of orthogonal projections. We give a method to reduce inequalities involving orthogonal projections and continuous functions on a Hilbert space to the finite dimensional case.
In Section 3.3, we give a new proof of Tsirelson’s inequality via this reduction method.
3.1 Basic properties of finite dimensional approximation
The definition of finite dimensional approximation of orthogonal projections is given below.
Definition 3.1.
Let
  1. (S1)
    $V$ be a separable Hilbert space,
  2. (S2)
    $P_{1},\dots,P_{m}$ be a finite family of orthogonal projections on $V$.
We say that a sequence $\left\{\left(V_{j},\left\{P_{j,1},\dots,P_{j,m}\right\}\right)\right\}_{j=1}^{\infty}$ is a finite dimensional approximation of the pair $\left(V,\left\{P_{1},\dots,P_{m}\right\}\right)$ if and only if the following conditions hold:
  1. (i)
    For any $j\in\mathbb{N}$, $V_{j}$ is a finite dimensional subspace of $V$ and $\{P_{j,k}\}_{k=1}^{m}$ is a finite family of orthogonal projections on $V_{j}$.
  2. (ii)
    For any $v\in V$, it follows that
    $\lim_{M\to\infty}\max\{\|v-\pi_{M}v\|,\|P_{1}v-P_{M,1}\pi_{M}v\|,\dots,\|P_{m}v-P_{M,m}\pi_{M}v\|\}=0,$
    where $\pi_{M}$ is the orthogonal projection from $V$ onto $V_{M}$.
We present two equivalent characterizations of finite dimensional approximation of orthogonal projections in separable Hilbert spaces.
Proposition 3.2.
Assume the setups $(S1),(S2)$ in Definition 3.1, and let
  1. (S3)
    $\{(V_{j},\{P_{j,1},\dots,P_{j,m}\})\}_{j=1}^{\infty}$ be a sequence such that $V_{j}$ is a finite dimensional subspace of $V$ and $P_{j,1},\dots,P_{j,m}$ are orthogonal projections on $V_{j}$ for any $j\in\mathbb{N}$,
then the following conditions are equivalent:
  1. (i)
    The sequence $\{(V_{j},\{P_{j,1},\dots,P_{j,m}\})\}_{j=1}^{\infty}$ is a finite dimensional approximation of $(V,\{P_{1},\dots,P_{m}\})$.
  2. (ii)
    For some dense countable subset $D=\{v_{i}\}_{i=1}^{\infty}$ of $V$ and any $i\in\mathbb{N}$,
    $\lim_{M\to\infty}\max\{\|v_{i}-\pi_{M}v_{i}\|,\|P_{1}v_{i}-P_{M,1}\pi_{M}v_{i}\|,\dots,\|P_{m}v_{i}-P_{M,m}\pi_{M}v_{i}\|\}=0,$
    where $\pi_{M}$ is the orthogonal projection from $V$ onto $V_{M}$.
  3. (iii)
    For any dense countable subset $D=\{v_{i}\}_{i=1}^{\infty}$ of $V$ and any $i\in\mathbb{N}$,
    $\lim_{M\to\infty}\max\{\|v_{i}-\pi_{M}v_{i}\|,\|P_{1}v_{i}-P_{M,1}\pi_{M}v_{i}\|,\dots,\|P_{m}v_{i}-P_{M,m}\pi_{M}v_{i}\|\}=0,$
    where $\pi_{M}$ is the orthogonal projection from $V$ onto $V_{M}$.
Proof.
It is clear that (i) $\implies$ (iii) $\implies$ (ii). We prove that (ii) $\implies$ (i). Let $v\in V$ and $\varepsilon>0$. There exist $i\in\mathbb{N}$ and $n_{0}\in\mathbb{N}$ such that for any $M\in[n_{0},\infty)\cap\mathbb{N}$
$\max\{\|v-v_{i}\|,\|v_{i}-\pi_{M}v_{i}\|,\|P_{1}v_{i}-P_{M,1}\pi_{M}v_{i}\|,\dots,\|P_{m}v_{i}-P_{M,m}\pi_{M}v_{i}\|\}<\dfrac{\varepsilon}{4}.$
Let $M\in[n_{0},\infty)\cap\mathbb{N}$ and $k\in\{1,\dots,m\}$. Since the operator norm of every orthogonal projection is at most $1$, we have
$\|v-\pi_{M}v\|=\|v_{i}-\pi_{M}v_{i}+(\mathrm{id}_{V}-\pi_{M})(v-v_{i})\|<\varepsilon,$
$\|P_{k}v-P_{M,k}\pi_{M}v\|=\|P_{k}v_{i}-P_{M,k}\pi_{M}v_{i}+(P_{k}-P_{M,k}\pi_{M})(v-v_{i})\|<\varepsilon.$
For any finite family of orthogonal projections on a separable Hilbert space $V$, there exists a finite dimensional approximation. This theorem will be applied in Section 3.2.
Theorem 3.3.
Assume the setups $(S1),(S2)$ in Definition 3.1. Then there exists a finite dimensional approximation of $(V,\{P_{1},\dots,P_{m}\})$.
Proof.
Let $D=\{v_{i}\}_{i=1}^{\infty}$ be a dense countable subset of $V$. Let $n\in\mathbb{N}$. For $i\in\{1,\dots,m\}$, define $W_{i}$ by $W_{i}=\langle P_{i}v_{1},\dots,P_{i}v_{n}\rangle$ and let $Q_{i}$ be the orthogonal projection onto $W_{i}$. Let $i\in\{1,\dots,m\}$ and $j\in\{1,\dots,n\}$. Since
$v_{j}=P_{i}v_{j}+(\mathrm{id}_{V}-P_{i})v_{j},P_{i}v_{j}\in W_{i},(\mathrm{id}_{V}-P_{i})v_{j}\perp W_{i},$
we have $P_{i}v_{j}=Q_{i}v_{j}$.
Define $W$ by $W=\sum_{i=1}^{m}W_{i}$. It follows that $W$ is a finite dimensional invariant subspace under $Q_{1},\dots,Q_{m}$. Therefore, there exists an orthonormal basis $\{x_{k}\}_{k=1}^{\dim W}$ of $W$ and an orthonormal basis $\{y_{l}\}_{l=1}^{\infty}$ of $W^{\perp}$. Since $V=W\oplus W^{\perp}$, there exists a sufficiently large $L\in\mathbb{N}$ such that $\|v_{j}-\pi_{L}v_{j}\|<\dfrac{1}{n}\quad(j=1,\dots,n)$, where $\pi_{L}$ is the orthogonal projection onto $\langle\{x_{k}\}_{k=1}^{\dim W}\cup\{y_{l}\}_{l=1}^{L}\rangle$. Fix such $L$ and define $L(n)=L$. Let $V_{L}$ be $\langle\{x_{k}\}_{k=1}^{\dim W}\cup\{y_{l}\}_{l=1}^{L}\rangle$. Since $Q_{i}|W^{\perp}=0$ for any $i\in\{1,\dots,m\}$, $V_{L}$ is invariant under $Q_{1},\dots,Q_{m}$.
Define $P_{L,1},\dots,P_{L,m}$ by $P_{L,i}=Q_{i}|V_{L}\quad(i=1,\dots,m)$. Then $V_{L}$ is an invariant subspace under $P_{L,1},\dots,P_{L,m}$. Hence $P_{L,1},\dots,P_{L,m}$ are orthogonal projections on $V_{L}$. Moreover, for any $i\in\{1,\dots,m\}$, $j\in\{1,\dots,n\}$
$\|P_{i}v_{j}-P_{L,i}\pi_{L}v_{j}\|=\|Q_{i}(v_{j}-\pi_{L}v_{j})\|<\dfrac{1}{n}.$
We conclude that $\{(V_{L(n)},\{P_{L(n),1},\dots,P_{L(n),m}\})\}_{n=1}^{\infty}$ is a finite dimensional approximation of $(V,\{P_{1},\dots,P_{m}\})$. ∎
Every finite dimensional approximation of orthogonal projections also approximates every polynomial of the projections, as follows. This proposition will also be applied in Section 3.2.
Proposition 3.4.
Assume the setups $(S1),(S2)$ in Definition 3.1, and let
  1. (S3)
    $\{(V_{j},\{P_{j,1},\dots,P_{j,m}\})\}_{j=1}^{\infty}$ be a finite dimensional approximation of
    $(V,\{P_{1},\dots,P_{m}\})$, and
  2. (S4)
    $f(x_{1},\dots,x_{m})\in\mathbb{C}\langle x_{1},\dots,x_{m}\rangle$.
Then, for any $v\in V$, we have
$\lim_{M\to\infty}\|f(P_{1},\dots,P_{m})v-f(P_{M,1},\dots,P_{M,m})\pi_{M}v\|=0.$
Proof.
Let $M=\deg(f)$. Define the complex numbers $a_{i_{1},\dots,i_{k}}$
$(k\in\{0,\dots,\deg(f)\},i_{1},\dots,i_{k}\in\{1,\dots,m\})$ by
$f(x_{1},\dots,x_{m})=\sum_{\begin{subarray}{c}k\in\{0,\dots,\deg(f)\},\\ i_{1},\dots,i_{k}\in\{1,\dots,m\}\end{subarray}}a_{i_{1},\dots,i_{k}}x_{i_{1}}\cdots x_{i_{k}}.$
Define $a,B$ by
$a$ $=\max\{|a_{i_{1},\dots,i_{k}}|\ |k\in\{0,\dots,\deg(f)\},i_{1},\dots,i_{k}\in\{1,\dots,m\}\},$
$B$ $=(a+1)(m+1)^{M+1}.$
Let $v\in V$ and $\varepsilon>0$. Then
$\left\{P_{i_{1}}\cdots P_{i_{k}}v\ \middle|k\in\{0,\dots,M\},i_{1},\dots,i_{k}\in\{1,\dots,m\}\right\}$
is a finite family of vectors. Hence, for any sufficiently large $n\in\mathbb{N}$ and any $k\in\{0,\dots,M\}$ and any $i_{1},\dots,i_{k}\in\{1,\dots,m\}$, we have
$\left\|P_{i_{1}}\cdots P_{i_{k}}v-P_{n,i_{1}}\cdots P_{n,i_{k}}\pi_{n}v\right\|<\dfrac{1}{B}\varepsilon.$
Therefore, for any sufficiently large $n\in\mathbb{N}$
$\left\|f(P_{1},\dots,P_{m})v-f(P_{n,1},\dots,P_{n,m})\pi_{n}v\right\|<\varepsilon.$
The following proposition presents a method for constructing finite dimensional approximations of orthogonal projections $P_{1},\dots,P_{m}$ on a separable Hilbert space $V$ via a decomposition into invariant subspaces under these projections. This proposition will be applied in Section 4.
Proposition 3.5.
Assume the setups $(S1),(S2)$ in Definition 3.1, and let
  1. (S3)
    $V=\bigoplus_{i\in I}W_{i}$ be an orthogonal direct sum decomposition, where $I$ is an at most countable set, and $W_{i}$ is an invariant subspace under $P_{1},\dots,P_{m}$ for any $i\in I$,
  2. (A1)
    $\{(W_{i,n},\{P_{i,n,1},\dots,P_{i,n,m}\})\}_{n=1}^{\infty}$ be a finite dimensional approximation of $(W_{i},\{P_{1}|W_{i},\dots,P_{m}|W_{i}\})$ for $i\in I$, and
  3. (S4)
    $\{I_{i}\}_{i=1}^{\infty}$ be a family of finite subsets of $I$ such that $I_{i}\subset I_{i+1}$ for any $i\in\mathbb{N}$ and $\bigcup_{i=1}^{\infty}I_{i}=I$,
then
$\left\{\left(\bigoplus_{i\in I_{n}}W_{i,n},\left\{\bigoplus_{i\in I_{n}}P_{i,n,1},\dots,\bigoplus_{i\in I_{n}}P_{i,n,m}\right\}\right)\right\}_{n=1}^{\infty}$
is a finite dimensional approximation of $(V,\{P_{1},\dots,P_{m}\})$.
Proof.
Let $v\in V$ and $\varepsilon>0$. Then there exists $n_{1}\in\mathbb{N}$ such that for any $n\in[n_{1},\infty)\cap\mathbb{N}$
$\left\|v-\sum_{j\in I_{n}}w_{j}\right\|<\dfrac{\varepsilon}{8},$
where $w_{j}=\pi_{W_{j}}v$ and $\pi_{W_{j}}$ is the orthogonal projection onto $W_{j}$ for $j\in I_{n}$. For $n\in\mathbb{N}$, let $\pi_{n}$ be $\bigoplus_{j\in I_{n}}\pi_{W_{j}}$.
Since $I_{n_{1}}$ is a finite set, there exists $n_{2}\in\mathbb{N}$ such that
$\left\|P_{j,k}w_{j}-P_{j,n,k}\pi_{j,n}w_{j}\right\|<\dfrac{\varepsilon}{4\#I_{n_{1}}}$
for any $j\in I_{n_{1}}$, $k\in\{1,\dots,m\}$, and $n\in[n_{2},\infty)\cap\mathbb{N}$. Here $\pi_{j,n}$ is the orthogonal projection onto $W_{j,n}$. For $n,n^{\prime}\in\mathbb{N}$, let $\pi_{n,n^{\prime}}$ be $\bigoplus_{j\in I_{n}}\pi_{W_{j,n^{\prime}}}$.
Let $k\in\{1,\dots,m\}$ and $n\in[\max\{n_{1},n_{2}\},\infty)\cap\mathbb{N}$. From the definition of $n_{1}$, we have $\|v-\pi_{n}v\|<\dfrac{\varepsilon}{8}$. It follows that $\|\pi_{n}v-\pi_{n_{1}}v\|<\dfrac{\varepsilon}{4}$. From the definition of $n_{2}$, we have $\left\|\bigoplus_{j\in I_{n_{1}}}P_{j,k}\pi_{n_{1}}v-\bigoplus_{j\in I_{n_{1}}}P_{j,n,k}\pi_{n_{1},n}v\right\|<\dfrac{\varepsilon}{4}$. We conclude that
$\left\|P_{k}v-\bigoplus_{j\in I_{n}}P_{j,n,k}\pi_{n,n}v\right\|<\varepsilon.$
It is possible to construct a finite dimensional approximation of a certain type of orthogonal projections on Hilbert tensor products, as follows. This proposition will be applied in Section 3.2.
Proposition 3.6.
Let
  1. (S1)
    $V_{1},\dots,V_{m}$ be a finite family of separable Hilbert spaces,
  2. (S2)
    $P_{i,1},\dots,P_{i,n_{i}}$ be a finite family of orthogonal projections on $V_{i}$ for $i\in\{1,\dots,m\}$, and
  3. (S3)
    $\{(V_{i,n},\{P_{i,n,1},\dots,P_{i,n,n_{i}}\})\}_{n=1}^{\infty}$ be a finite dimensional approximation of $(V_{i},\{P_{i,1},\dots,P_{i,n_{i}}\})$ for $i\in\{1,\dots,m\}$,
then
$\left\{\left(\bigotimes_{l=1}^{m}V_{l,n},\left\{P_{i,n,k_{i}}\middle|k_{i}\in[1,n_{i}]\cap\mathbb{N}\quad(i=1,\dots,m)\right\}\right)\right\}_{n=1}^{\infty}$
is a finite dimensional approximation of
$\left(\bigotimes_{l=1}^{m}V_{l},\left\{P_{i,k_{i}}\middle|k_{i}\in[1,n_{i}]\cap\mathbb{N}\quad(i=1,\dots,m)\right\}\right).$
Proof.
Let $\psi\in\bigotimes_{i=1}^{m}V_{i}$ and $\varepsilon>0$. From the definition of $\bigotimes_{i=1}^{m}V_{i}$, the linear span $\langle\{v_{1}\otimes\dots\otimes v_{m}|v_{i}\in V_{i}\quad(i=1,\dots,m)\}\rangle$ is dense in $\bigotimes_{i=1}^{m}V_{i}$. Therefore, there exist $M\in\mathbb{N}$ and $v_{i,1},\dots,v_{i,M}\in V_{i}$ $(i=1,\dots,m)$ such that $\left\|\psi-\sum_{k=1}^{M}v_{1,k}\otimes\dots\otimes v_{m,k}\right\|<\dfrac{\varepsilon}{3}$. Define $L$ by
$L=\max\left\{\|v_{i,k}\|\ |i=1,\dots,m,k=1,\dots,M\right\}+1.$
There exists $n_{0}\in\mathbb{N}$ such that for any $n\in[n_{0},\infty)\cap\mathbb{N}$, $i\in\{1,\dots,m\}$, $j\in\{1,\dots,n_{i}\}$, and $k\in\{1,\dots,M\}$, we have
$\left\|P_{i,j}v_{i,k}-P_{i,n,j}\pi_{i,n}v_{i,k}\right\|<\dfrac{\varepsilon}{3ML^{m}}.$
Here, $\pi_{i,n}$ is the orthogonal projection onto $V_{i,n}$ for $i\in\{1,\dots,m\}$ and $n\in\mathbb{N}$. Let $n\in[n_{0},\infty)\cap\mathbb{N}$. Then, we have
$\left\|P_{i,j_{i}}\sum_{k=1}^{M}\bigotimes_{l=1}^{m}v_{l,k}-P_{i,n,j_{i}}\bigotimes_{l=1}^{m}\pi_{l,n}\sum_{k=1}^{M}\bigotimes_{l=1}^{m}v_{l,k}\right\|<\dfrac{\varepsilon}{3}$
for any $i\in\{1,\dots,m\}$, $j_{i}\in\{1,\dots,n_{i}\}$. Note that $\bigotimes_{l=1}^{m}\pi_{l,n}$ is the orthogonal projection onto $\bigotimes_{l=1}^{m}V_{l,n}$. We conclude that for any $i\in\{1,\dots,m\}$, $j_{i}\in\{1,\dots,n_{i}\}$,
$\left\|P_{i,j_{i}}\psi-P_{i,n,j_{i}}\bigotimes_{l=1}^{m}\pi_{l,n}\psi\right\|<\varepsilon.$
For a polynomial $f(P_{1},\dots,P_{m})$ of orthogonal projections $P_{1},\dots,P_{m}$ on a Hilbert space $V$, we define the finite dimensional approximate point spectrum of $f(P_{1},\dots,P_{m})$ corresponding to a finite dimensional approximation $\{(V_{n},\{P_{n,1},\dots,P_{n,m}\})\}_{n=1}^{\infty}$, as follows. Such a spectrum consists of all complex numbers $\lambda$ satisfying the following conditions:
  • 1.
    $\lambda$ is approximated by eigenvalues $\lambda_{n}$ $(n\in\mathbb{N})$ of $f(P_{n,1},\dots,P_{n,m})$.
  • 2.
    The actions of $P_{1},\dots,P_{m}$ on an eigenvector corresponding to $\lambda_{n}$ are approximated by $P_{n,1},\dots,P_{n,m}$.
This notion plays a role in Section 5.
Definition 3.7.
Assume the setups $(S1),(S2)$ in Definition 3.1, and let
  1. (S3)
    $f\in\mathbb{C}\langle x_{1},\dots,x_{m}\rangle$,
  2. (S4)
    $\{(V_{n},\{P_{n,1},\dots,P_{n,m}\})\}_{n=1}^{\infty}$ be a finite dimensional approximation of $(V,\{P_{1},\dots,P_{m}\})$.
We say $\lambda\in\mathbb{C}$ is in the finite dimensional approximate point spectrum of $f(P_{1},\dots,P_{m})$ corresponding to $\{(V_{n},\{P_{n,1},\dots,P_{n,m}\})\}_{n=1}^{\infty}$ if there exists a sequence $\{\lambda_{n}\}_{n=1}^{\infty}$ in $\mathbb{C}$ and a sequence $\{u_{n}\}_{n=1}^{\infty}$ of unit vectors in $V$ such that
  1. (i)
    For any $n\in\mathbb{N}$, $\lambda_{n}$ is an eigenvalue of $f(P_{n,1},\dots,P_{n,m})$.
  2. (ii)
    For any $n\in\mathbb{N}$, $u_{n}$ is an eigenvector of $f(P_{n,1},\dots,P_{n,m})$ corresponding to $\lambda_{n}$.
  3. (iii)
    $\lim_{n\to\infty}\lambda_{n}=\lambda$.
  4. (iv)
    For any $j\in\{1,\dots,m\}$, $\lim_{n\to\infty}\|P_{j}u_{n}-P_{n,j}u_{n}\|=0$.
We note that every complex number in the finite dimensional approximate point spectrum of $f(P_{1},\dots,P_{m})$ lies in the approximate point spectrum of $f(P_{1},\dots,P_{m})$. For the definition of the approximate point spectrum, see, e.g., [17].
3.2 Reduction of inequalities to the finite dimensional case
In various fields, including quantum mechanics, one encounters the following problem: for a Hilbert space $V$, a non-commutative polynomial $f$ of orthogonal projections $P_{1},\dots,P_{m}$, and a real-valued continuous function $F$ on $V$, estimate the following quantity,
$\sup_{\begin{subarray}{c}\Psi\in V,\|\Psi\|=1\end{subarray}}F(f(P_{1},\dots,P_{m})\Psi),$
where $F$ represents the quantity to be estimated.
In quantum information theory, one often considers problems where $V$ is a Hilbert tensor product $\bigotimes_{i=1}^{m}V_{i}$, and orthogonal projections are $\operatorname{id}_{V_{1}}\otimes\dots\otimes P_{i,j}\otimes\dots\otimes\operatorname{id}_{V_{m}}$ ($i\in\{1,\dots,m\}$, $j\in\{1,\dots,n\}$).
As an example, consider the problem of estimating the spectral radius of the Bell-CHSH operator. In this problem, we have $m=2$, and $f$ and $F$ are given as follows.
Example 3.8.
Let
  1. (S1)
    $V_{1}$, $V_{2}$ be separable Hilbert spaces,
  2. (S2)
    $P_{i,j}$ be an orthogonal projection on $V_{i}$ for $i,j\in\{1,2\}$,
  3. (S3)
    the non-commutative polynomial $f$ be defined by
    $f(x_{1,1},\dots,x_{2,2})=2(x_{1,1}+x_{1,2}-1)(2x_{2,1}-1)+2(x_{1,1}-x_{1,2})(2x_{2,2}-1),\text{ and}$
  4. (S4)
    the real-valued continuous function $F$ on $V_{1}\otimes V_{2}$ be defined by
    $F(\Psi)=\|\Psi\|\quad(\Psi\in V_{1}\otimes V_{2}),$
then the following identity holds:
$\sup_{\Psi\in V_{1}\otimes V_{2},\|\Psi\|=1}F(f(P_{1,1},P_{1,2},P_{2,1},P_{2,2})\Psi)=\rho(\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})).$
Proof.
Define $A_{1},A_{2},B_{1},B_{2}$ by $A_{i}=2P_{1,i}-\operatorname{id}_{V_{1}}$, $B_{i}=2P_{2,i}-\operatorname{id}_{V_{2}}$ $(i=1,2)$. Then
$\quad\mathcal{B}=(A_{1}+A_{2})\otimes B_{1}+(A_{1}-A_{2})\otimes B_{2}$
$=2(P_{1,1}+P_{1,2}-\operatorname{id}_{V_{1}})\otimes(2P_{2,1}-\operatorname{id}_{V_{2}})+2(P_{1,1}-P_{1,2})\otimes(2P_{2,2}-\operatorname{id}_{V_{2}})$
$=f(P_{1,1},P_{1,2},P_{2,1},P_{2,2}).$
Since $\mathcal{B}$ is a self-adjoint operator, its spectral radius coincides with its operator norm. Therefore, we have
$\quad\rho(\mathcal{B})=\|\mathcal{B}\|=\sup_{\begin{subarray}{c}\Psi\in V_{1}\otimes V_{2},\\ \|\Psi\|=1\end{subarray}}\|\mathcal{B}\Psi\|=\sup_{\begin{subarray}{c}\Psi\in V_{1}\otimes V_{2},\\ \|\Psi\|=1\end{subarray}}F(f(P_{1,1},P_{1,2},P_{2,1},P_{2,2})\Psi).$
The following Theorem 3.9 allows us to reduce the above problem to the case where the target Hilbert spaces are finite dimensional. As an application, we present a new proof of Tsirelson’s inequality in Section 3.3. Another application is a security proof of device-independent quantum key distribution against a class of attacks known as collective attacks, which will be presented in our forthcoming work [18].
Theorem 3.9.
Assume the setups $(S1),(S2)$ in Definition 3.1, and let
  1. (S3)
    $F$ be a real-valued continuous function on $V$,
  2. (S4)
    $f(x_{1},\dots,x_{m})\in\mathbb{C}\langle x_{1},\dots,x_{m}\rangle$, and
  3. (A1)
    there exists a constant $C\in\mathbb{R}$ such that for any finite dimensional subspace $V_{f}$ of $V$, any orthogonal projections $P_{f,1},\dots,P_{f,m}$ on $V_{f}$, and any unit vector $\Psi_{f}\in V_{f}$, it follows that
    $F\left(f\left(P_{f,1},\dots,P_{f,m}\right)\Psi_{f}\right)\leq C,$
then for any unit vector $\Psi\in\bigotimes_{i=1}^{m}V_{i}$, we have
$F\left(f\left(P_{1},\dots,P_{m}\right)\Psi\right)\leq C.$
Proof.
From Theorem 3.3, there exists a finite dimensional approximation $\{(V_{n},\{P_{n,1},\dots,P_{n,m}\})\}_{n=1}^{\infty}$ of $(V,\{P_{1},\dots,P_{m}\})$. Let $\Psi$ be a unit vector in $V$. From Proposition 3.4, we have
$\lim_{n\to\infty}\left\|f\left(P_{1},\dots,P_{m}\right)\Psi-f\left(P_{n,1},\dots,P_{n,m}\right)\dfrac{\pi_{n}\Psi}{\|\pi_{n}\Psi\|}\right\|=0,$
where $\pi_{n}$ is the orthogonal projection onto $V_{n}$. From assumption (A1), for any $n\in\mathbb{N}$, it follows that
$F\left(f\left(P_{n,1},\dots,P_{n,m}\right)\dfrac{\pi_{n}\Psi}{\|\pi_{n}\Psi\|}\right)\leq C.$
Since $F$ is continuous, we conclude that
$F\left(f\left(P_{1},\dots,P_{m}\right)\Psi\right)\leq C.$
From Theorem 3.9 and Proposition 3.6, we have the following corollary.
Corollary 3.10.
Let
  1. (S1)
    $V_{1},\dots,V_{m}$ be a finite family of separable Hilbert spaces,
  2. (S2)
    $P_{i,1},\dots,P_{i,n}$ be a finite family of orthogonal projections on $V_{i}$ for $i\in\{1,\dots,m\}$,
  3. (S3)
    $F$ be a real-valued continuous function on $\bigotimes_{i=1}^{m}V_{i}$,
  4. (S4)
    $f(x_{1,1},\dots,x_{i,j},\dots,x_{m,n})\in\mathbb{C}\langle x_{1,1},\dots,x_{i,j},\dots,x_{m,n}\rangle$, and
  5. (A1)
    there exists a constant $C\in\mathbb{R}$ such that for any $i\in\{1,\dots,m\}$, any finite dimensional subspace $V_{f,i}$ of $V_{i}$, any orthogonal projections $P_{f,i,1},\dots,P_{f,i,n}$ on $V_{f,i}$, and any unit vector $\Psi_{f}\in\bigotimes_{i=1}^{m}V_{f,i}$, it follows that
    $F\left(f\left(P_{f,1,1},\dots,P_{f,i,j},\dots,P_{f,m,n}\right)\Psi_{f}\right)\leq C,$
then for any unit vector $\Psi\in\bigotimes_{i=1}^{m}V_{i}$, we have
$F\left(f\left(P_{1,1},\dots,P_{i,j},\dots,P_{m,n}\right)\Psi\right)\leq C.$
3.3 Application: New proof of Tsirelson’s inequality
First, note that Tsirelson’s inequality (2.2) is straightforward to prove in the case where either subsystem of the bipartite quantum system has dimension at most two.
Lemma 3.11.
We assume the setups and conditions of Theorem 2.4, and suppose that $\dim V_{1}\leq 2$. Then inequality (2.2) holds.
Proof.
If $\dim V_{1}=1$, (2.2) follows by simple calculus. Thus, we can assume that $\dim V_{1}=2$ and $\tau_{P_{1,1},P_{1,2}}$ is irreducible. From Fact 2.2 and Fact 2.3, there exists $x\in(0,1)$ such that
$\mathcal{B}(P_{1,1},P_{1,2},P_{2,1},P_{2,2})$
$=\begin{pmatrix}2x&2\sqrt{x(1-x)}\\ 2\sqrt{x(1-x)}&-2x\end{pmatrix}\otimes B_{1}+\begin{pmatrix}-2(1-x)&2\sqrt{x(1-x)}\\ 2\sqrt{x(1-x)}&2(1-x)\end{pmatrix}\otimes B_{2}.$
Define $C,D$ by
$C=\begin{pmatrix}2x&2\sqrt{x(1-x)}\\ 2\sqrt{x(1-x)}&-2x\end{pmatrix},D=\begin{pmatrix}-2(1-x)&2\sqrt{x(1-x)}\\ 2\sqrt{x(1-x)}&2(1-x)\end{pmatrix}$.
From Fact 2.3, the eigenvalues of $C$ are $2\sqrt{x},2\sqrt{x}$ and the eigenvalues of $D$ are $2\sqrt{1-x},2\sqrt{1-x}$. Therefore, the spectral radius $\rho(\mathcal{B})$ is at most $2\sqrt{x}+2\sqrt{(1-x)}$. From simple differential calculus, we conclude that
$0,1]\ni x\mapsto 2\sqrt{x}+2\sqrt{(1-x)}\in\mathbb{R}$
takes the maximum value at $x=1/2$ and the maximum value is $2\sqrt{2}$. Therefore, inequality (2.2) holds. ∎
Our new proof of Tsirelson’s inequality is the following.
New proof of Tsirelson’s inequality.
From Example 3.8 and Corollary 3.10, we may assume that $\dim V_{1}<\infty$. Moreover, from Fact 2.2, there exists a finite family $\{W_{i}\}_{i=1}^{m}$ of invariant subspaces under $P_{1,1}$ and $P_{1,2}$ such that $\dim W_{i}\leq 2$ for any $i\in\{1,\dots,m\}$ and $V_{1}=\bigoplus_{i=1}^{m}W_{i}$ and
$W_{i_{1}}\perp W_{i_{2}}\quad(\text{for any }i_{1},i_{2}\in\{1,\dots,m\}\text{ such that }i_{1}\neq i_{2}).$
Since
  • 1.
    $W_{i}\otimes V_{2}$ is invariant under $\mathcal{B}$ for any $i\in\{1,\dots,m\}$,
  • 2.
    $V_{1}\otimes V_{2}=\bigoplus_{i=1}^{m}W_{i}\otimes V_{2}$,
  • 3.
    $W_{i_{1}}\otimes V_{2}\perp W_{i_{2}}\otimes V_{2}\quad$ for any $i_{1},i_{2}\in\{1,\dots,m\}$ such that $i_{1}\neq i_{2}$,
we may assume that $\dim V_{1}\leq 2$. From Lemma 3.11, we conclude that inequality (2.2) holds. ∎
We remark on the difference between our proof and others. There are many proofs of Tsirelson’s inequality. Several of such proofs are reviewed in [19]. The proof based on (2.1) is known to be particularly simple.
Tsirelson [13] established an ingenious technique using the Clifford algebra to reduce Tsirelson’s inequality to the case where $\dim V_{i}\leq 2$ $(i=1,2)$ in Theorem 2.4. Our proof presents another approach that reduces Tsirelson’s inequality to the two dimensional case, using Corollary 3.10 and Fact $2.$
4  One-shifted form of two orthogonal projections
We introduce one-shifted form matrix representations for pairs of orthogonal projections on infinite dimensional Hilbert spaces, and present their basic properties and applications. Let $V$ be a separable infinite dimensional Hilbert space, and $P,Q$ be orthogonal projections on $V$.
In Section 4.1, we define one-shifted form matrix representations and present their basic properties. In particular, we prove that if $P$ and $Q$ admit a one-shifted form then they admit an explicit finite dimensional approximation.
In Section 4.2, we present a complete orthogonal direct sum decomposition of $V$ into invariant subspaces $\{V_{i}\}_{i\in I}$ under $P$ and $Q$ such that for any $i\in I$ $\dim(V_{i})\leq 2$, or $P|V_{i}$ and $Q|V_{i}$ admit a one-shifted form. This decomposition is an extension of Jordan’s lemma for finite dimensional Hilbert spaces.
In Section 4.3, we present an explicit formula for finite dimensional approximation of pairs of orthogonal projections, based on the extension of Jordan’s lemma.
4.1 Basic properties of one-shifted form
We define one-shifted form matrix representations for pairs of orthogonal projections as follows.
Definition 4.1 (One-shifted form).
Let
  1. (S1)
    $V$ be an infinite dimensional Hilbert space,
  2. (S2)
    $P$ and $Q$ be orthogonal projections on $V$, and
  3. (S3)
    $\mathcal{U}=\{u_{1},v_{1},u_{2},v_{2},\dots\}$ be an orthonormal basis of $V$.
We say that the pair $(V,\{P,Q\})$ admits a one-shifted form in $\mathcal{U}$ if and only if there are two sequences $\Theta=\{\theta_{i}\}_{i=1}^{\infty},\Omega=\{\omega_{i}\}_{i=1}^{\infty}$ of angles in $(0,\pi)$ such that the following conditions hold:
  1. (i)
    For any $i\in\mathbb{N}$, the spanned subspace $\langle u_{i},v_{i}\rangle$ is invariant under $P$. The matrix representation of $2P|\langle u_{i},v_{i}\rangle-\operatorname{id}_{\langle u_{i},v_{i}\rangle}$ is
    $\begin{pmatrix}\cos\theta_{i}&\sin\theta_{i}\\ \sin\theta_{i}&-\cos\theta_{i}\end{pmatrix}$.
  2. (ii)
    The equality $Qu_{1}=u_{1}$ holds and $\langle v_{i},u_{i+1}\rangle$ is invariant under $Q$ for any $i\in\mathbb{N}$. The matrix representation of $2Q|\langle v_{i},u_{i+1}\rangle-\operatorname{id}_{\langle v_{i},u_{i+1}\rangle}$ is
    $\begin{pmatrix}\cos\omega_{i}&\sin\omega_{i}\\ \sin\omega_{i}&-\cos\omega_{i}\end{pmatrix}$.
In particular, $V$ is separable. The matrix representation of $2P-\operatorname{id}_{V}$ is
[…]
and the matrix representation of $2Q-\operatorname{id}_{V}$ is
[…]
We present two characterizations of pairs of orthogonal projections that admit a one-shifted form. The second characterization (ii) via cyclic vectors is employed in an extension of Jordan’s lemma, which is discussed in Theorem 4.4. The third characterization (iii) via explicit approximations is employed to derive explicit formulas for finite dimensional approximations of pairs of orthogonal projections, which is stated in Theorem 4.6.
Proposition 4.2.
Assume the setups $(S1),(S2)$ in Definition 4.1, and let
  1. (A1)
    $\dim V=\infty$,
then the following three conditions are equivalent.
  1. (i)
    The pair $(V,\{P,Q\})$ admits a one-shifted form.
  2. (ii)
    There exists a cyclic unit vector $u\in V$ for $\tau_{P,Q}$ such that $Qu=u$.
  3. (iii)
    There exists an orthonormal basis $\mathcal{U}=\{u_{1},v_{1},u_{2},v_{2},\dots\}$ of $V$ and sequences $\Theta=(\theta_{i})_{i=1}^{\infty},\Omega=(\omega_{i})_{i=1}^{\infty}$ of angles in $(0,\pi)$ such that
    $\{(\langle u_{1},v_{1},\dots,u_{n},v_{n}\rangle,\{P_{n}(\Theta),Q_{n}(\Omega)\})\}_{n=1}^{\infty}$ is a finite dimensional approximation of $(V,\{P,Q\})$.
Proof of that (i) $\implies$ (ii) and (iii) $\implies$ (i):.
These implications are straightforward. ∎
Proof of that (i) $\implies$ (iii):.
Let $\mathcal{U}=\{u_{1},v_{1},u_{2},v_{2},\dots\}$ be an orthonormal basis of $V$ such that $(V,\{P,Q\})$ has a one-shifted form in $\mathcal{U}$. Let $v\in V$ and $\varepsilon>0$. For a sufficiently large $n\in\mathbb{N}$, there exists $w\in\langle u_{1},v_{1},\dots,u_{n},v_{n}\rangle$ such that $\|v-w\|<\varepsilon$. Then $Pw=P_{n+1}(\Theta)w$ and $Qw=Q_{n+1}(\Omega)w$. Hence, condition (iii) is satisfied. ∎
Step 1 in the proof of (ii) $\implies$ (i):.
In this step, we construct $u_{1},v_{1},u_{2}$. Define $D$ and $u_{1}$ by $D=\{v\in V|\ \|v\|=1\},u_{1}=u$. If $Pu_{1}\in\langle u_{1}\rangle$, $\langle u_{1}\rangle$ is an invariant subspace under $\tau_{P,Q}$. Since $u$ is a cyclic vector, we have $\langle u_{1}\rangle=V$. That contradicts assumption (A1). Therefore, the vector $Pu_{1}$ is not in $\langle u_{1}\rangle$. Hence, there exists $v_{1}\in D$ such that $v_{1}\in\langle u_{1},Pu_{1}\rangle$ and $v_{1}\perp u_{1}$ and $\langle u_{1},v_{1}\rangle$ is invariant under $P$. For any $\phi_{1}\in\mathbb{R}$, $\langle u_{1},\exp(i\phi_{1})v_{1}\rangle$ is also invariant under $P$ and $P$ is positive semidefinite on $\langle u_{1},\exp(i\phi_{1})v_{1}\rangle$. Therefore, we may assume that the matrix representation of $P$ on $\langle u_{1},\exp(i\phi_{1})v_{1}\rangle$ is $\begin{pmatrix}x&y\\ y&z\end{pmatrix}$ for some $x,y,z\in[0,\infty)$. Since $\begin{pmatrix}x&y\\ y&z\end{pmatrix}^{2}=\begin{pmatrix}x&y\\ y&z\end{pmatrix}$, we have $x^{2}+y^{2}=x,xy+yz=y$. Therefore, the equalities $y=\sqrt{x(1-x)},z=1-x$ hold. It follows that there exists $\psi_{1}\in\left(0,\dfrac{\pi}{2}\right)$ such that $\sqrt{x}=\cos\psi_{1},\sqrt{1-x}=\sin\psi_{1}$. Therefore, we have $\begin{pmatrix}x&y\\ y&z\end{pmatrix}=\begin{pmatrix}\cos^{2}\psi_{1}&\cos\psi_{1}\sin\psi_{1}\\ \cos\psi_{1}\sin\psi_{1}&1-\cos^{2}\psi_{1}\end{pmatrix}$. Moreover, it follows that
$\begin{pmatrix}\cos^{2}\psi_{1}&\cos\psi_{1}\sin\psi_{1}\\ \cos\psi_{1}\sin\psi_{1}&1-\cos^{2}\psi_{1}\end{pmatrix}$ $=\dfrac{1}{2}\begin{pmatrix}\cos(2\psi_{1})+1&\sin(2\psi_{1})\\ \sin(2\psi_{1})&1-\cos(2\psi_{1})\end{pmatrix}$
$=\dfrac{1}{2}\begin{pmatrix}\cos\theta_{1}&\sin\theta_{1}\\ \sin\theta_{1}&-\cos\theta_{1}\end{pmatrix}+\dfrac{1}{2}E_{2},$
where $\theta_{1}=2\psi_{1}\in(0,\pi)$.
If $Qv_{1}\in\langle u_{1},v_{1}\rangle$, $\langle u_{1},v_{1}\rangle$ is an invariant subspace under $\tau_{P,Q}$. Since $u$ is a cyclic vector, we have $\langle u_{1},v_{1}\rangle=V$. It follows that $\dim V\leq 2$. That contradicts assumption (A1). Therefore, we have $Qv_{1}\not\in\langle u_{1},v_{1}\rangle$. It follows that there exists $u_{2}\in D$ such that $u_{2}\in\langle v_{1},Qv_{1}\rangle$ and $u_{2}\perp v_{1}$. Then $\langle v_{1},u_{2}\rangle$ is invariant under $Q$. Since $u_{2}=av_{1}+bQv_{1}$ for some $a,b\in\mathbb{C}$, we have $(u_{2},u_{1})=a(v_{1},u_{1})+b(Qv_{1},u_{1})=b(v_{1},Qu_{1})=0$. Therefore, we conclude that $u_{2}\perp\langle u_{1},v_{1}\rangle$. Following the same reasoning as above, we may assume that there exists $\omega_{1}\in(0,\pi)$ such that the matrix representation of $Q|\langle v_{1},u_{2}\rangle$ is
$\dfrac{1}{2}\begin{pmatrix}\cos\omega_{1}&\sin\omega_{1}\\ \sin\omega_{1}&-\cos\omega_{1}\end{pmatrix}+\dfrac{1}{2}E_{2}.$
Consequently, the vectors $u_{1},v_{1}$ and $u_{2}$ lie in $D$, are mutually orthogonal, and satisfy the following conditions:
  • 1.
    The spanned subspace $\langle u_{1},v_{1}\rangle$ is invariant under $P$, and the matrix representation of $P|\langle u_{1},v_{1}\rangle$ is
    $\dfrac{1}{2}\begin{pmatrix}\cos\theta_{1}&\sin\theta_{1}\\ \sin\theta_{1}&-\cos\theta_{1}\end{pmatrix}+\dfrac{1}{2}E_{2}\quad(\theta_{1}\in(0,\pi)).$
  • 2.
    The equality $Qu_{1}=u_{1}$ holds, and $\langle v_{1},u_{2}\rangle$ is invariant under $Q$, and the matrix representation of $Q|\langle v_{1},u_{2}\rangle$ is
    $\dfrac{1}{2}\begin{pmatrix}\cos\omega_{1}&\sin\omega_{1}\\ \sin\omega_{1}&-\cos\omega_{1}\end{pmatrix}+\dfrac{1}{2}E_{2}\quad(\omega_{1}\in(0,\pi)).$
Step 2 in the proof of (ii) $\implies$ (i):.
In this step, we construct the orthogonal basis $\{u_{1},v_{1},u_{2},v_{2},\dots\}$ by induction. Let $i\in\mathbb{N}$ and assume that we get $\{u_{1},\dots,u_{i},v_{i},u_{i+1}\}$ satisfy the claim in condition (i). Following the same reasoning as Step 1, we have $Pu_{i+1}\not\in\langle u_{1},\dots,u_{i},v_{i},u_{i+1}\rangle$. Hence, there exists $v_{i+1}\in D$ such that $v_{i+1}\in\langle u_{i+1},Pu_{i+1}\rangle$ and $v_{i+1}\perp u_{i+1}$ and $\langle u_{i+1},v_{i+1}\rangle$ is invariant under $P$. As in Step 1, we may assume that the matrix representation of $P|\langle u_{i+1},v_{i+1}\rangle$ is
$\dfrac{1}{2}\begin{pmatrix}\cos\theta_{i+1}&\sin\theta_{i+1}\\ \sin\theta_{i+1}&-\cos\theta_{i+1}\end{pmatrix}+\dfrac{1}{2}E_{2}\quad\text{ for some $\theta_{i+1}\in(0,\pi)$}.$
Let $w\in\langle u_{1},\dots,u_{i},v_{i}\rangle$. Since $v_{i+1}=au_{i+1}+bPu_{i+1}$ for some $a,b\in\mathbb{C}$ and $\langle u_{1},\dots,u_{i},v_{i}\rangle$ is invariant under $P$, we have $(v_{i+1},w)=a(u_{i+1},w)+b(Pu_{i+1},w)=b(u_{i+1},Pw)=0$. Therefore, we have
$v_{i+1}\perp\langle u_{1},v_{1},\dots,u_{i},v_{i},u_{i+1}\rangle.$
As in Step 1, it follows that $Qv_{i+1}\not\in\langle u_{1},v_{1},\dots,u_{i+1},v_{i+1}\rangle$. Hence, there exists $u_{i+2}\in D$ such that $u_{i+2}\in\langle v_{i+1},Qv_{i+1}\rangle$ and $u_{i+2}\perp v_{i+1}$. Let $w\in\langle u_{1},v_{1},\dots,u_{i+1}\rangle$. Since $u_{i+2}=av_{i+1}+bQv_{i+1}$ for some $a,b\in\mathbb{C}$ and $\langle u_{1},v_{1},\dots,u_{i+1}\rangle$ is invariant under $Q$, we have $(u_{i+2},w)=a(v_{i+1},w)+b(Qv_{i+1},w)=b(v_{i+1},Qw)=0$. Therefore, we have
$u_{i+2}\perp\langle u_{1},v_{1},\dots,u_{i+1},v_{i+1}\rangle$
and $\langle v_{i+1},u_{i+2}\rangle$ is invariant under $Q$.
Following the same reasoning as Step 1, we may assume that there exists $\omega_{i+1}\in(0,\pi)$ such that the matrix representation of $Q|\langle v_{i+1},u_{i+2}\rangle$ is
$\dfrac{1}{2}\begin{pmatrix}\cos\omega_{i+1}&\sin\omega_{i+1}\\ \sin\omega_{i+1}&-\cos\omega_{i+1}\end{pmatrix}+\dfrac{1}{2}E_{2}.$
Thus we get $u_{1},\dots,u_{i+1},v_{i+1},u_{i+2}$ satisfying the claim in condition (i).
By induction, we get $\mathcal{U}=\{u_{1},v_{1},\dots\}\subset D$ such that
  • 1.
    For any $i\in\mathbb{N}$, the spanned subspace $\langle u_{i},v_{i}\rangle$ is invariant under $P$ and the matrix representation of $P|\langle u_{i},v_{i}\rangle$ is
    $\dfrac{1}{2}\begin{pmatrix}\cos\theta_{i}&\sin\theta_{i}\\ \sin\theta_{i}&-\cos\theta_{i}\end{pmatrix}+\dfrac{1}{2}E_{2}\quad(\theta_{i}\in(0,\pi)).$
  • 2.
    The equality $Qu_{1}=u_{1}$ holds, and for any $i\in\mathbb{N}$ $\langle v_{i},u_{i+1}\rangle$ is invariant under $Q$ and the matrix representation of $Q|\langle v_{i},u_{i+1}\rangle$ is
    $\dfrac{1}{2}\begin{pmatrix}\cos\omega_{i}&\sin\omega_{i}\\ \sin\omega_{i}&-\cos\omega_{i}\end{pmatrix}+\dfrac{1}{2}E_{2}\quad(\omega_{i}\in(0,\pi)).$
Moreover, $\langle u_{1},v_{1},\dots,u_{i},v_{i},\dots\rangle$ is invariant under $\tau_{P,Q}$. From the assumption that $u$ is a cyclic vector, the set $\mathcal{U}$ is an orthonormal basis of $V$. ∎
We introduce the following notation for constant-angle one-shifted forms.
Notation 4.3.
Let $\theta\in(0,\pi)$ and define $\Theta=\{\theta_{i}\}_{i=1}^{\infty}$ and $\Omega=\{\omega_{i}\}_{i=1}^{\infty}$ by $\theta_{i}=\omega_{i}=\theta$ $(i=1,2,\dots)$. Define self-adjoint operators $A(\theta),B(\theta)$ on $\ell^{2}$ by $A(\theta)=A(\Theta)$ and $B(\theta)=B(\Omega)$. We identify $\mathbb{C}^{n}$ with
$\left\{\{a_{i}\}_{i=1}^{n}\in\ell^{2}\middle|a_{n+k}=0\quad(k=1,2,\dots)\right\}$. Define the finite dimensional approximation $\mathcal{A}$ of $(\ell^{2},\{P(\theta),Q(\theta)\})$ by $\mathcal{A}=\{(\mathbb{C}^{n},\{P_{n}(\theta),Q_{n}(\theta)\})\}_{n=1}^{\infty}$, where $P_{n}(\theta)=P_{n}(\Theta)$, $Q_{n}(\theta)=Q_{n}(\Omega)$, $A_{n}(\theta)=2P_{n}(\theta)-E_{2n}$, $B_{n}(\theta)=2Q_{n}(\theta)-E_{2n}$ $(n=1,2,\dots)$.
4.2 Extension of Jordan’s lemma
An extension of Jordan’s lemma is stated in Theorem 4.4. From the viewpoint of group representation theory, Theorem 4.4 provides an explicit formula that decomposes every infinite dimensional unitary representation of $D_{\infty}$ into cyclic representations.
Theorem 4.4.
Assume the setups $(S1),(S2)$ in Definition 4.1, then there exist at most countable closed invariant subspaces $\{W_{i}\}_{i\in I}$ under $\tau_{P,Q}$ such that
  1. (i)
    The orthogonal direct sum decomposition $V=\bigoplus_{i\in I}W_{i}$ holds.
  2. (ii)
    For any $i\in I$, the subspace $W_{i}$ is at most two dimensional or infinite dimensional.
  3. (iii)
    For any $i\in I$ such that $\dim W_{i}\leq 2$, $W_{i}$ is irreducible under $\tau_{P,Q}$.
  4. (iv)
    For any $i\in I$ such that $\dim W_{i}=\infty$, the pair $(W_{i},\{P|W_{i},Q|W_{i}\})$ has a one-shifted form and $W_{i}$ does not admit nonzero finite dimensional invariant subspaces under $\tau_{P,Q}$.
Proof.
We define the sets $\mathcal{B},\mathcal{C},\mathcal{D},\mathcal{T}$ by
[…]
and admits no nonzero finite dimensional invariant subspaces under $\tau_{P,Q}$
[…]
$\mathcal{D}$ $=\mathcal{B}\cup\mathcal{C},$
$\mathcal{T}$ $=\left\{D\subset\mathcal{D}\middle|W_{1}\perp W_{2}\quad(\text{for any $W_{1},W_{2}\in D$ such that }W_{1}\neq W_{2})\right\}.$
It is straightforward that $\mathcal{T}$ is not empty. We assume that the order of $\mathcal{T}$ is defined by set inclusion. Let $T$ be a totally ordered subset of $\mathcal{T}$. It follows directly that $\bigcup_{D\in T}D\in\mathcal{T}$. From Zorn’s Lemma, the set $\mathcal{T}$ has a maximal element $D$. We define $V_{0}$ by $V_{0}=\bigoplus_{W\in D}W.$ To derive a contradiction, we assume $V_{0}\neq V$. If $Q|V_{0}^{\perp}=0$ or $V_{0}^{\perp}$ admits a nonzero finite dimensional invariant subspace of $V$ under $\tau_{P,Q}$, from Fact 2.2, the subspace $V_{0}^{\perp}$ has an at most two dimensional invariant subspace $W$ under $\tau_{P,Q}$. Then $W\in\mathcal{B}$, and hence $D\cup\{W\}\in\mathcal{T}$. That contradicts the fact that $D$ is a maximal element of $\mathcal{T}$.
Assume $Q|V_{0}^{\perp}\neq 0$ and $V_{0}^{\perp}$ admits no nonzero finite dimensional invariant subspaces of $V$ under $\tau_{P,Q}$. Then there exists a unit vector $u\in V_{0}^{\perp}$ such that $Qu=u$. Define $W$ by $W=\overline{\langle\tau_{P,Q}(D_{\infty})u\rangle}$. It is straightforward that $W\perp V_{0}$. Since $V_{0}^{\perp}$ does not admit nonzero finite dimensional invariant subspaces of $V$ under $\tau_{P,Q}$, the subspace $W$ also does not admit such subspaces and $\dim W=\infty$. Then $W\in\mathcal{C}$ and $W\perp W^{\prime}$ for any $W^{\prime}\in D$. Therefore, we have $D\cup\{W\}\in\mathcal{T}$. That contradicts the fact that $D$ is a maximal element of $\mathcal{T}$. Therefore, we have $V=\bigoplus_{W\in D}W$. Since $V$ is separable, $D$ must be at most countable. From Proposition 4.2, for any $W\in\mathcal{C}$, $(W,\{P|W,Q|W\})$ has a one-shifted form. ∎
Note that there exists a pair of orthogonal projections $P$ and $Q$ on an infinite dimensional Hilbert space $V$ satisfying the following conditions:
  • 1.
    The pair $(V,\{P,Q\})$ admits a constant-angle one-shifted form.
  • 2.
    There does not exist a nonzero finite dimensional invariant subspace under $P$ and $Q$.
Example 4.5.
Let
  1. (S1)
    $V,V_{1},V_{2}$ be defined by
    $V$ $=\ell^{2}=\left\{\{x_{j}\}_{j=1}^{\infty}\subset\mathbb{C}\middle|\sum_{j=1}^{\infty}|x_{j}|^{2}<\infty\right\},$
    $V_{1}$ $=\left\{\{x_{j}\}_{j=1}^{\infty}\in\ell^{2}\middle|x_{2k+1}=x_{2k+2}\quad(k=0,1,2,\dots)\right\},$
    $V_{2}$ $=\left\{\{x_{j}\}_{j=1}^{\infty}\in\ell^{2}\middle|x_{2k+2}=x_{2k+3}\quad(k=0,1,2,\dots)\right\},$
  2. (S2)
    $P_{i}$ for $i\in\{1,2\}$ be the orthogonal projection onto $V_{i}$ and $A_{i}$ for $i\in\{1,2\}$ be defined by $A_{i}=2P_{i}-\operatorname{id}_{V}$,
then
  1. (i)
    Let $x=\{x_{j}\}_{j=1}^{\infty}\in V$. Then the following holds:
    $P_{1}x$ $=(\dfrac{1}{2}(x_{1}+x_{2}),\dfrac{1}{2}(x_{1}+x_{2}),\dots,$
    $\quad\dfrac{1}{2}(x_{2k+1}+x_{2k+2}),\dfrac{1}{2}(x_{2k+1}+x_{2k+2}),\dots),$
    $A_{1}x$ $=(x_{2},x_{1},\dots,x_{2k+2},x_{2k+1},\dots),$
    $P_{2}x$ $=(x_{1},\dfrac{1}{2}(x_{2}+x_{3}),\dfrac{1}{2}(x_{2}+x_{3}),\dots,$
    $\quad\dfrac{1}{2}(x_{2k+2}+x_{2k+3}),\dfrac{1}{2}(x_{2k+2}+x_{2k+3}),\dots),$
    $A_{2}x$ $=(x_{1},x_{3},x_{2},\dots,x_{2k+3},x_{2k+2},\dots).$
  2. (ii)
    Let $x=\{x_{j}\}_{j=1}^{\infty}\in V$ and $k\in\mathbb{N}$. Then
    $(A_{1}A_{2}x)_{4k+1}$ $=x_{4k+3},$ $(A_{1}A_{2}x)_{4k+3}$ $=x_{4k+5},$
    $(A_{1}A_{2}x)_{4k+2}$ $=x_{4k},$ $(A_{1}A_{2}x)_{4k+4}$ $=x_{4k+2},$
    $(A_{2}A_{1}x)_{4k+1}$ $=x_{4k-1},$ $(A_{2}A_{1}x)_{4k+3}$ $=x_{4k+1},$
    $(A_{2}A_{1}x)_{4k+2}$ $=x_{4k+4},$ $(A_{2}A_{1}x)_{4k+4}$ $=x_{4k+6}.$
  3. (iii)
    The pair $(V,\{P_{1},P_{2}\})$ admits a constant-angle one-shifted form whose angle is $\dfrac{\pi}{2}$ in the standard basis.
  4. (iv)
    There does not exist a nonzero finite dimensional invariant subspace under $P_{1}$ and $P_{2}$.
Proof of (i)-(iii).
These follow directly from the definitions of $P_{1}$, $P_{2}$, $A_{1}$, and $A_{2}$, and straightforward calculations. ∎
Proof of (iv).
Let $W$ be a nonzero invariant subspace under $\tau_{P_{1},P_{2}}$. From Fact 2.2, it is enough to show that $\dim W\geq 3$. Since $W$ is nonzero, there exists $w\in W\setminus\{0\}$. From statement (ii), there are $m\in\mathbb{N}$ and a sequence $i_{1},\dots,i_{m}\in\{1,2\}$ and $\alpha\in\mathbb{C}$ such that the first component of $\alpha A_{i_{1}}\dots A_{i_{m}}w$ is $1$. Without loss of generality, we may assume that $w_{1}=1$. Since $w\in\ell^{2}$, we have $\lim_{n\to\infty}w_{n}=0$. Therefore, there exists $i_{0}\in[2,\infty)\cap\mathbb{N}$ such that $|w_{i_{0}}|\neq 1$ and $w_{j}=1$ for any $j\in\{1,\dots,i_{0}-1\}$. From statement (ii), there exists a finite sequence $j_{1},\dots,j_{n}\in\{1,2\}$ such that $A_{j_{1}}\dots A_{j_{n}}w=(1,w^{\prime}_{2},\dots),A_{1}A_{j_{1}}\dots A_{j_{n}}w=(w^{\prime}_{2},1,\dots),$ where $|w^{\prime}_{2}|\neq 1$. Since $(w^{\prime}_{2})^{2}\neq 1$, these two vectors are linearly independent. Hence, there exist $u,v\in W$ such that $u=(1,0,u_{3},\dots),v=(0,1,v_{3},\dots)$.
To derive a contradiction, assume that $\dim W=2$. Since $A_{2}u=(1,u_{3},\dots)$ is in $\langle u,v\rangle$, we have $u_{3}=0$. Since $A_{2}v=(0,v_{3},\dots)$ is in $\langle u,v\rangle$, we have $|v_{3}|=1$. Therefore, it follows that $u=(1,0,0,u_{4},\dots),v=(0,1,v_{3},v_{4},\dots)$ and $|v_{3}|=1$. Since $A_{1}u=(0,1,u_{4},0,\dots)$ is in $\langle u,v\rangle$, we have $|u_{4}|=1$. Since $A_{1}v=(1,0,v_{4},v_{3},\dots)$ is in $\langle u,v\rangle$, we have $v_{4}=0$. Therefore, the following holds for $k=0$:
$|u_{4k+1}|$ $=1,$ $u_{4k+2}$ $=0,$ $u_{4k+3}$ $=0,$ $|u_{4k+4}|$ $=1,$
$v_{4k+1}$ $=0,$ $|v_{4k+2}|$ $=1,$ $|v_{4k+3}|$ $=1,$ $v_{4k+4}$ $=0.$
Let $k_{0}\in[0,\infty)\cap\mathbb{Z}$ and assume that the above statement holds for $k=k_{0}$. From statement (ii), the following equalities hold for some $\{a_{i}\}_{i=1}^{8}\subset\mathbb{C}$:
$(A_{1}A_{2})^{k_{0}+1}u$ $=(0,a_{1},u_{4(k_{0}+1)+1},\dots),$
$(A_{1}A_{2})^{k_{0}+1}v$ $=(v_{4k_{0}+3},a_{2},v_{4(k_{0}+1)+1},\dots),$
$(A_{1}A_{2})^{k_{0}+1}A_{1}u$ $=(u_{4(k_{0}+1)},a_{3},u_{4(k_{0}+1)+2},\dots),$
$(A_{1}A_{2})^{k_{0}+1}A_{1}v$ $=(0,a_{4},v_{4(k_{0}+1)+2},\dots),$
$(A_{1}A_{2})^{k_{0}+2}u$ $=(u_{4(k_{0}+1)+1},a_{5},u_{4(k_{0}+1)+3},\dots),$
$(A_{1}A_{2})^{k_{0}+2}v$ $=(v_{4(k_{0}+1)+1},a_{6},v_{4(k_{0}+1)+3},\dots),$
$(A_{1}A_{2})^{k_{0}+2}A_{1}u$ $=(u_{4(k_{0}+1)+2},a_{7},u_{4(k_{0}+1)+4},\dots),$
$(A_{1}A_{2})^{k_{0}+2}A_{1}v$ $=(v_{4(k_{0}+1)+2},a_{8},v_{4(k_{0}+1)+4},\dots),$
and these vectors are in $\langle u,v\rangle$. Hence, the following equalities hold:
$|u_{4(k_{0}+1)+1}|$ $=1,$ $v_{4(k_{0}+1)+1}$ $=0$
$u_{4(k_{0}+1)+2}$ $=0,$ $|v_{4(k_{0}+1)+2}|$ $=1$
$u_{4(k_{0}+1)+3}$ $=0,$ $|v_{4(k_{0}+1)+3}|$ $=1$
$|u_{4(k_{0}+1)+4}|$ $=1,$ $v_{4(k_{0}+1)+4}$ $=0.$
By induction, the above statement holds for any $k\in[0,\infty)\cap\mathbb{Z}$. This conclusion contradicts the fact that $\lim_{k\to\infty}u_{k}=0$. ∎
4.3 Explicit construction of finite dimensional approximations
From Theorem 4.4 and Proposition 3.5, we obtain an explicit formula for finite dimensional approximation of pairs of orthogonal projections, as follows.
Theorem 4.6.
Assume the setting of Theorem 4.4, and let
  1. (S3)
    $\{I_{n}\}_{n=1}^{\infty}$ be a sequence of finite subsets of $I$ such that $I_{n}\subset I_{n+1}$ for any $n\in\mathbb{N}$ and $\bigcup_{n=1}^{\infty}I_{n}=I$,
  2. (S4)
    $\mathcal{U}=\{u_{i,1},v_{i,1},u_{i,2},v_{i,2},\dots\}$ be an orthonormal basis of $W_{i}$, and $\Theta_{i}=\{\theta_{i,j}\}_{j=1}^{\infty}$ and $\Omega_{i}=\{\omega_{i,j}\}_{j=1}^{\infty}$ be in $(0,\pi)$ for $i\in I$ such that $\dim W_{i}=\infty$,
  3. (S5)
    $(W_{i},P|W_{i},Q|W_{i})$ admits a one-shifted form in $\mathcal{U}$ whose angles are $\Theta_{i},\Omega_{i}$ for any $i\in I$ such that $\dim W_{i}=\infty$,
  4. (S6)
    for $n\in\mathbb{N}$ and $i\in I_{n}$, $W_{i,n}$ be defined by
    $W_{i,n}=\left\{\begin{array}[]{ll}\langle u_{i,1},v_{i,1},\dots,u_{i,n},v_{i,n}\rangle&\quad\dim W_{i}=\infty,\\ W_{i}&\quad\text{otherwise},\end{array}\right$.
  5. (S7)
    for $n\in\mathbb{N}$ and $i\in I_{n}$, $P_{i,n}$ and $Q_{i,n}$ be defined by
    $P_{i,n}=\left\{\begin{array}[]{ll}P_{n}(\Theta_{i})&\quad\dim W_{i}=\infty,\\ P|W_{i}&\quad\text{otherwise},\end{array}\right.\quad Q_{i,n}=\left\{\begin{array}[]{ll}Q_{n}(\Omega_{i})&\quad\dim W_{i}=\infty,\\ Q|W_{i}&\quad\text{otherwise},\end{array}\right$.
    and
  6. (S8)
    $V_{n},P_{n},Q_{n}$ for $n\in\mathbb{N}$ be defined by
    $V_{n}=\bigoplus_{i\in I_{n}}W_{i,n},P_{n}=\bigoplus_{i\in I_{n}}P_{i,n},Q_{n}=\bigoplus_{i\in I_{n}}Q_{i,n},$
then the sequence $\{(V_{n},\{P_{n},Q_{n}\})\}_{n=1}^{\infty}$ is a finite dimensional approximation of $(V,\{P,Q\})$.
5  Spectral radius of commutators of two orthogonal projections
We present a new method for estimating the spectral radius $\rho([P,Q])=\rho([A,B])$ for a pair of orthogonal projections $P$ and $Q$ on a Hilbert space $V$, where $A=2P-\operatorname{id}_{V}$ and $B=2Q-\operatorname{id}_{V}$. As stated in Section 2.3, this problem is equivalent to estimating the spectral radius of the Bell-CHSH operator. Moreover, from Theorem 4.4, we can reduce the problem of estimating $\rho([A,B])$ to the following two cases:
  1. Case 1
    $\tau_{P,Q}$ admits an orthogonal direct sum decomposition into finite dimensional invariant subspaces under $\tau_{P,Q}$,
  2. Case 2
    $\tau_{P,Q}$ does not satisfy Case 1 and admits a one-shifted form.
Case 1 is trivial. Hence we mainly consider Case 2.
In Section 5.1, we give upper and lower estimates of $\rho([A,B])$ in Case 1 and Case 2.
In Section 5.2, we prove that these estimates become exact when $P$ and $Q$ admit a constant-angle one-shifted form.
In Section 5.3, we give upper and lower estimates of the spectral radius of the Bell-CHSH operator in the case where every subsystem of the bipartite quantum system satisfies Case 1 or Case 2.
5.1 Upper and lower estimates of the spectral radius
In this section, we use the following notation.
Notation 5.1.
Let
  1. (S1)
    $V$ be a separable Hilbert space,
  2. (S2)
    $P$ and $Q$ be orthogonal projections on $V$,
  3. (S3)
    AA and $B$ be defined by $A=2P-\operatorname{id}_{V},B=2Q-\operatorname{id}_{V}$, and
  4. (S4)
    $\mathcal{A}=\{(V_{n},\{P_{n},Q_{n}\})\}_{n=1}^{\infty}$ be a finite dimensional approximation of
    $(V,\{P,Q\})$.
With this notation, let $F(A+B,\mathcal{A})$ be the set of all real numbers $\lambda\in(0,2)$ such that $\lambda$ and $\lambda$ are in the finite dimensional approximate point spectrum of $A+B=2(P+Q-\operatorname{id}_{V})$ corresponding to $\mathcal{A}$. Moreover, let $\tilde{F}(A+B,\mathcal{A})$ be defined by
$\tilde{F}(A+B,\mathcal{A})=F(A+B,\mathcal{A})\cup\{0\}.$
We denote $F(A+B,\mathcal{A}),\tilde{F}(A+B,\mathcal{A})$ by $F(A+B),\tilde{F}(A+B)$, respectively, whenever no confusion arises.
In Theorem 5.3, which is the main result of this section, we provide upper and lower bounds for $\rho([A,B])$. The upper bound is established in the general setting, without assuming Case 1 or Case 2. The lower bound is derived under the assumption that either Case 1 or Case 2 holds. As a preparation for Theorem 5.2, we first identify $F(A+B)$ in Case 1. In Case 1, Fact 2.3 allows us to directly identify set $F(A+B)$, as follows.
Lemma 5.2.
Assume the setups (S1)-(S3) in Notation 5.1, and that
  1. (A1)
    there exists an at most countable family $\{W_{i}\}_{i\in I}$ of invariant subspaces under $P$ and $Q$ such that
    $V=\bigoplus_{\begin{subarray}{c}i\in I\end{subarray}}W_{i},W_{i_{1}}\perp W_{i_{2}}\ (\text{for any $i_{1}\neq i_{2}\in I$}),\text{ and let }$
  2. (S4)
    $\{I_{n}\}_{n=1}^{\infty}$ be a family of finite subsets of $I$ such that $I_{n}\subset I_{n+1}$ for any $n\in\mathbb{N}$ and $\bigcup_{n=1}^{\infty}I_{n}=I$,
  3. (S5)
    $\mathcal{A}$ be a finite dimensional approximation of $(V,\{P,Q\})$ defined by
    $\mathcal{A}=\{(V_{n},\{P_{n},Q_{n}\})\}_{n=1}^{\infty},V_{n}=\bigoplus_{i\in I_{n}}W_{i},P_{n}=\bigoplus_{i\in I_{n}}P|W_{i},Q_{n}=\bigoplus_{i\in I_{n}}Q|W_{i},$
then we have
$F(A+B,\mathcal{A})=2\overline{\bigcup_{i\in I}\sigma((P+Q)|W_{i}-\operatorname{id}_{W_{i}})}\cap(0,2).$
We give upper and lower estimates of the spectral radius of the commutator $[A,B]$, as follows.
Theorem 5.3.
Assume the setups in Notation 5.1, and let
  1. (S5)
    $A_{n}$ and $B_{n}$ be defined by $A_{n}=2P_{n}-\operatorname{id}_{V_{n}}$, $B_{n}=2Q_{n}-\operatorname{id}_{V_{n}}$ for $n\in\mathbb{N}$,
then the following statements hold:
  1. (i)
    Let $\lambda_{n}$ be the eigenvalue of $A_{n}+B_{n}$ whose absolute value is closest to $\sqrt{2}$ for $n\in\mathbb{N}$. Then
    $\rho([A,B])\leq\liminf_{n\to\infty}b(\lambda_{n}).$
  2. (ii)
    Assume that $(V,\{P,Q\})$ has a one-shifted form and $\mathcal{A}$ ( in (S4) of Notation 5.1 ) is the finite dimensional approximation of $(V,\{P,Q\})$ defined in Theorem 4.6. Then
    $\sup_{\lambda\in\tilde{F}(A+B,\mathcal{A})}b(\lambda)\leq\rho([A,B]).$
  3. (iii)
    Assume that $(V,\{P,Q\})$ satisfies the conditions (S1)-(S3), (A1) in Lemma 5.2. Then
    $\sup_{\lambda\in\tilde{F}(A+B,\mathcal{A})}b(\lambda)=\rho([A,B])=\liminf_{n\to\infty}b(\lambda_{n}).$
Proof of (i):.
Let $\{(V_{\phi(j)},\{P_{\phi(j)},Q_{\phi(j)}\})\}_{j=1}^{\infty}$ be a subsequence of $\mathcal{A}$ such that $\{b(\lambda_{\phi(j)})\}_{j=1}^{\infty}$ converges, and let $v\in V$ and $\varepsilon>0$. From Proposition 3.4, there exists $n_{0}\in\mathbb{N}$ such that for any $n\in[n_{0},\infty)\cap\mathbb{N}$, we have $|([A,B]v,v)|-\varepsilon<|([A_{n},B_{n}]\pi_{n}v,\pi_{n}v)|$. From Fact 2.3, for any $n\in[n_{0},\infty)\cap\mathbb{N}$
$|([A_{n},B_{n}]\pi_{n}v,\pi_{n}v)|\leq\rho([A_{n},B_{n}])=b(\lambda_{n}).$
Hence, we have $|([A,B]v,v)|-\varepsilon\leq\liminf_{n\to\infty}b(\lambda_{n})$. Since $\varepsilon$ is arbitrary, we conclude that $|([A,B]v,v)|\leq\liminf_{n\to\infty}b(\lambda_{n})$. ∎
Proof of (ii).
If $F(A+B)=\emptyset$, it is straightforward that $\sup_{\lambda\in\tilde{F}(A+B,\mathcal{A})}b(\lambda)\leq\rho([A,B])$. Thus we may assume that $F(A+B)\neq\emptyset$. Let $\lambda\in F(A+B)$. Let $\{\lambda_{i}\}_{i=1}^{\infty}$ be a sequence in $(0,2)$, and let $\{u_{i}\}_{i=1}^{\infty}$ and $\{v_{i}\}_{i=1}^{\infty}$ be sequences of unit vectors in $V$ such that
  • 1.
    For any $i\in\mathbb{N}$, $\lambda_{i}$ and $\lambda_{i}$ are eigenvalues of $A_{i}+B_{i}$.
  • 2.
    For any $i\in\mathbb{N}$, $u_{i}$ and $v_{i}$ are eigenvectors of $A_{i}+B_{i}$ corresponding to $\lambda_{i}$ and $\lambda_{i}$, respectively.
  • 3.
    The following convergence holds:
    $\lim_{i\to\infty}\lambda_{i}=\lambda,\lim_{i\to\infty}\max\{\|(A-A_{i})u_{i}\|,\|(A-A_{i})v_{i}\|\}=0.$
Let $i\in\mathbb{N}$. Then all subdiagonal elements of $A_{i}+B_{i}$ are nonzero. From Fact 2.5, the eigenvalues of $A_{i}+B_{i}$ are distinct. From Fact 2.2 and Fact 2.3, there exists a two dimensional irreducible invariant subspace $W$ under $\tau_{P_{i},Q_{i}}$ such that $(A_{i}+B_{i})|W$ has eigenvalues $\lambda_{i},\lambda_{i}$ and $u_{i},v_{i}\in W$. From Fact 2.3, there exist $\alpha_{i},\beta_{i}\in\mathbb{C}$ such that $|\alpha_{i}|^{2}+|\beta_{i}|^{2}=1$ and
$|([A_{i},B_{i}]w_{i},w_{i})|=\rho([A_{i}|W,B_{i}|W])=b(\lambda_{i}),$
where $w_{i}=\alpha_{i}u_{i}+\beta_{i}v_{i}$. From the definition of the spectral radius of $[A,B]$, we have $|([A,B]w_{i},w_{i})|\leq\rho([A,B])$. Let $\varepsilon>0$. For any sufficiently large $i\in\mathbb{N}$, the inequality $|[A,B]w_{i}-[A_{i},B_{i}]w_{i}|<\varepsilon$ holds. Then
$b(\lambda_{i})-\varepsilon=|([A_{i},B_{i}]w_{i},w_{i})|-\varepsilon\leq|([A,B]w_{i},w_{i})|\leq\rho([A,B]).$
Since $b$ is continuous, by letting $i\to\infty$, we have $b(\lambda)-\varepsilon\leq\rho([A,B])$. Since $\varepsilon$ is arbitrary, we conclude that $b(\lambda)\leq\rho([A,B])$. ∎
Proof of (iii).
It follows from Lemma 5.2. ∎
5.2 Equality conditions for the spectral radius estimates
First, we prove some properties of the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ for $n\in\mathbb{N}$.
Lemma 5.4.
Given $\theta\in(0,\pi)$, let
  1. (S1)
    Functions $f_{n}:\mathbb{R}\setminus\left\{\dfrac{k}{n+1}\pi\middle|k\in\mathbb{Z}\right\}\to\mathbb{R}$ for $n\in\mathbb{N}$ be defined by
    $f_{n}(\phi)=\dfrac{\sin((n-1)\phi)}{\sin((n+1)\phi)}\quad(\phi\in\mathbb{R}),$
  2. (S2)
    Intervals $I_{n,k}$ for $n\in\mathbb{N}$ and $k\in[0,n]\cap\mathbb{Z}$ be defined by
    $I_{n,k}=\left(\dfrac{k}{n+1}\pi,\dfrac{k+1}{n+1}\pi\right),\text{ and}$
  3. (S3)
    Sets $Z_{n}$ for $n\in\mathbb{N}$ be defined by
    $Z_{n}=\left\{\phi\in\mathbb{R}\middle|(1+\cos\theta)^{2}\sin((n-1)\phi)=\sin^{2}(\theta)\sin((n+1)\phi)\right\},$
then the following statements hold:
  1. (i)
    For any $n\in\mathbb{N}$ and $\phi\in[0,\pi]$,
    $f_{2n}(\phi)=f_{2n}(\pi-\phi)=f_{2n}(\pi+\phi).$
  2. (ii)
    For any sufficiently large $n\in\mathbb{N}$, $f_{2n}$ is monotone in each of the following two intervals:
    $\left(0,\dfrac{\pi}{2(2n+1)}\right),\left(\pi-\dfrac{\pi}{2(2n+1)},\pi\right).$
  3. (iii)
    For any sufficiently large $n\in\mathbb{N}$ and $k\in([1,2n-1]\cap\mathbb{N})\setminus\{n\}$, $f_{2n}$ is monotone in each of $\left[\dfrac{\pi}{2(2n+1)},\dfrac{\pi}{2n+1}\right)$, $I_{2n,k}$,
    $\left(\dfrac{2n\pi}{2n+1},\pi-\dfrac{\pi}{2(2n+1)}\right]$.
  4. (iv)
    Assume that $\theta\neq\dfrac{1}{2}\pi$. Then, for any sufficiently large $n\in\mathbb{N}$,
    $\#(Z_{2n}\cap I_{2n,0})\leq 3,\#(Z_{2n}\cap I_{2n,2n})\leq 3,Z_{2n}\cap I_{2n,n}=\emptyset,$
    and $\{0,\pi\}\subset Z_{2n}$, and for any $k\in([1,2n-1]\cap\mathbb{N})\setminus\{n\}$
    $\#(Z_{2n}\cap I_{2n,k})=1.$
  5. (v)
    Assume that $\theta=\dfrac{1}{2}\pi$. Then
    $Z_{2n}\cap[0,\pi]\subset\left\{0,\pi\right\}\cup\left\{\dfrac{1}{4n}\pi,\dfrac{3}{4n}\pi,\dots,\dfrac{4n-1}{4n}\pi\right\}.$
  6. (vi)
    Assume that $\theta\neq\dfrac{1}{2}\pi$. Then for any sufficiently large $n\in\mathbb{N}$,
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,0})\sin\theta)\leq 3,$
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,2n})\sin\theta)\leq 3,$
    $\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,n})\sin\theta=\emptyset,$
    and for any $k\in([1,2n-1]\cap\mathbb{N})\setminus\{n\}$,
    $\#(\sigma(A_{n}(\theta)+B_{n}(\theta))\cap 2\cos(I_{2n,k})\sin\theta)\leq 1.$
  7. (vii)
    Assume that $\theta=\dfrac{1}{2}\pi$. Then for any $n\in\mathbb{N}$
    $\sigma(A_{n}(\theta)+B_{n}(\theta))\subset\{-2,2\}\cup\left\{2\cos\left(\dfrac{2k-1}{4n}\pi\right)\middle|k=1,2,\dots,2n\right\}.$
  8. (viii)
    Let $\varepsilon>0$ and $\lambda_{0}\in(-2\sin\theta,2\sin\theta)$ and $\eta>0$ such that $|\lambda_{0}|+\eta<2\sin\theta$. Then there exist $n_{0}\in\mathbb{N}$ such that for any $n\in[n_{0},\infty)\cap\mathbb{N}$ and any eigenvalue $\lambda$ of $A_{n}(\theta)+B_{n}(\theta)$ such that $\lambda\in(\lambda_{0}-\eta,\lambda_{0}+\eta)$, and for any unit eigenvector $u$ corresponding to $\lambda$, we have $|u_{2n}|<\varepsilon$.
Proof of (i).
From the fundamental trigonometric identity, we have
$\sin((2n+1)(\pi-\phi))$ $=\sin((2n+1)\phi),$
$\sin((2n-1)(\pi-\phi))$ $=\sin((2n-1)\phi),$
$\sin((2n+1)(\pi+\phi))$ $=(-1)\sin((2n+1)\phi),$
$\sin((2n-1)(\pi+\phi))$ $=(-1)\sin((2n-1)\phi).$
Thus, (i) holds. ∎
Proof of (ii).
Let $n\in\mathbb{N}$. From statement (i), it is enough to prove that $f_{2n}$ is monotone in $\left(0,\dfrac{\pi}{2(2n+1)}\right)$. Let $\phi\in\left(0,\dfrac{\pi}{2(2n+1)}\right)$. Then we have
$f_{2n}^{\prime}(\phi)$ $=\dfrac{1}{\sin^{2}((2n+1)\phi)}\{(2n-1)\cos((2n-1)\phi)\sin((2n+1)\phi)$
$\quad-(2n+1)\sin((2n-1)\phi)\cos((2n+1)\phi)\}$
$=\dfrac{(2n-1)\cos((2n-1)\phi)\sin((2n+1)\phi)}{\sin^{2}((2n+1)\phi)}\left\{1-\dfrac{b\tan(a\phi)}{a\tan(b\phi)}\right\},$
where a=2n1,b=2n+1a=2n-1,b=2n+1. There exists a sequence $\{c_{i}\}_{i=0}^{\infty}$ of positive numbers such that for any constant $C>0$ and $x\in\left[0,\dfrac{\pi}{2C}\right)$, $\tan(Cx)=\sum_{i=0}^{\infty}c_{i}C^{2i+1}x^{2i+1}$. Therefore, we have $1-\dfrac{b\tan(a\phi)}{a\tan(b\phi)}>0$. Thus $f_{2n}$ is monotone in $\left(0,\dfrac{\pi}{2(2n+1)}\right)$. ∎
Proof of (iii).
Let $\phi$ be a number in
$\left[\dfrac{\pi}{2(2n+1)},\dfrac{\pi}{2n+1}\right)\cup\bigcup_{k\in\{1,\dots,2n-1\}\setminus\{n\}}I_{2n,k}\cup\left(\dfrac{2n\pi}{2n+1},\pi-\dfrac{\pi}{2(2n+1)}\right].$
Then, the following equality holds:
$f_{2n}^{\prime}(\phi)$
$=\dfrac{1}{\sin^{2}((2n+1)\phi)}\{(2n-1)\cos((2n-1)\phi)\sin((2n+1)\phi)$
$\quad-(2n+1)\sin((2n-1)\phi)\cos((2n+1)\phi)\}$
$=\dfrac{(2n-1)}{\sin^{2}((2n+1)\phi)}\{\cos((2n-1)\phi)\sin((2n+1)\phi)$
$\quad-\sin((2n-1)\phi)\cos((2n+1)\phi)$
$\quad+\left(1-\dfrac{(2n+1)}{(2n-1)}\right)\sin((2n-1)\phi)\cos((2n+1)\phi)\}$
$=\dfrac{1}{\sin^{2}((2n+1)\phi)}\left\{(2n-1)\sin(2\phi)-2\sin((2n-1)\phi)\cos((2n+1)\phi)\right\}.$
Here,
$\dfrac{1}{2n+1}\pi\leq 2\phi\leq\dfrac{2n}{2n+1}\pi\quad\text{or}\quad\dfrac{2n+2}{2n+1}\pi\leq 2\phi\leq 2\pi-\dfrac{\pi}{2n+1}.$
Hence, we have
$|(2n-1)\sin(2\phi)|\geq\left(2n-1\right)\min\left\{a,b,c,d\right\},$
where $a,b,c$ are defined by
$a$ $=\left|\sin\left(\dfrac{1}{2n+1}\pi\right)\right|,b=\left|\sin\left(\dfrac{2n}{2n+1}\pi\right)\right|,$
$c$ $=\left|\sin\left(\dfrac{2n+2}{2n+1}\pi\right)\right|,d=\left|\sin\left(2\pi-\dfrac{\pi}{2n+1}\right)\right|$
From the proof of (i), we have
$\left|\sin\left(\dfrac{2n}{2n+1}\pi\right)\right|=\left|\sin\left(\pi-\dfrac{1}{2n+1}\pi\right)\right|=\left|\sin\left(\dfrac{1}{2n+1}\pi\right)\right|,$
$\left|\sin\left(\dfrac{2n+2}{2n+1}\pi\right)\right|=\left|\sin\left(\pi+\dfrac{1}{2n+1}\pi\right)\right|=\left|\sin\left(\dfrac{1}{2n+1}\pi\right)\right|.$
Since for any sufficiently large $n\in\mathbb{N}$
$\left(2n-1\right)\sin\left(\dfrac{1}{2n+1}\pi\right)=\sin\left(\dfrac{1}{2n+1}\pi\right)\dfrac{2n+1}{\pi}\dfrac{2n-1}{2n+1}\pi>2,$
we have $|(2n-1)\sin(2\phi)|>2$. Since $|2\sin((2n-1)\phi)\cos((2n+1)\phi)|\leq 2$, for any sufficiently large $n\in\mathbb{N}$, $f_{2n}$ is monotone in each of the following intervals:
$\left[\dfrac{\pi}{2(2n+1)},\dfrac{\pi}{2n+1}\right),I_{2n,1},\dots,I_{2n,n-1},I_{2n,n+1},\dots,I_{2n,2n-1},$
$\left(\dfrac{2n\pi}{2n+1},\pi-\dfrac{\pi}{2(2n+1)}\right].$
Proof of (iv).
Since $\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}}=\tan^{2}\left(\dfrac{\theta}{2}\right)$, we have $\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}}\neq 1$. For any sufficiently large $n\in\mathbb{N}$,
$\dfrac{2n-1}{2n+1}\neq\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}},\quad\sin\left(\dfrac{\pi}{2n+1}\right)\dfrac{2n+1}{\pi}\dfrac{2n-1}{2n+1}\pi>2,$
and for any $\phi\in I_{2n,n}$ $|f_{2n}(\phi)-1|<1-\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}}$.
Fix such an $n\in\mathbb{N}$. It follows that $Z_{2n}\cap I_{2n,n}=\emptyset$.
From statements (ii) and (iii), we have $(Z_{2n}\cap I_{2n,0}),\#(Z_{2n}\cap I_{2n,2n})\leq 3$.
Let $k\in\{1,\dots,2n-1\}\setminus\{n\}$. From statement (iii) and
$\lim_{\eta\to\frac{k}{2n+1}+0}|f_{2n}(\eta)|=\lim_{\eta\to\frac{k+1}{2n+1}-0}|f_{2n}(\eta)|=\infty,$
$\lim_{\begin{subarray}{c}\eta_{1}\to\frac{k}{2n+1}+0,\\ \eta_{2}\to\frac{k+1}{2n+1}-0\end{subarray}}f_{2n}(\eta_{1})f_{2n}(\eta_{2})=-\infty,$
we have $(Z_{2n}\cap I_{2n,k})=1$. ∎
Proof of (v).
Let $\phi\in Z_{2n}$. Since $\dfrac{\sin^{2}\theta}{(1+\cos\theta)^{2}}=1$, we have
$\sin((2n-1)\phi)=\sin((2n+1)\phi).$
Therefore, there exists $k\in\mathbb{Z}$ such that
$4n\phi=(2k+1)\pi\text{ or }2\phi=2k\pi.$
Hence, statement (v) holds. ∎
Proof of (vi)-(viii).
We will first prove (vi) and (vii). In order to do so, we apply Yuen’s result [16], which is introduced in Section 2.4.2, to the symmetric tridiagonal matrix
$A(\theta)+B(\theta)=\begin{pmatrix}1+\cos\theta&\sin\theta&0&\dots&0&0\\ \sin\theta&0&\sin\theta&\dots&0&0\\ 0&\sin\theta&0&\dots&0&0\\ \dots&\dots&\dots&\dots&\dots&\dots\\ 0&0&0&\dots&0&\sin\theta\\ 0&0&0&\dots&\sin\theta&-(1+\cos\theta)\end{pmatrix}$.
Let $a,b,c,\alpha,\beta,\gamma,\delta,\mu$ be complex numbers defined in Section 2.4.2. In the present setting,
$a=c=\sin\theta\neq 0,\quad\alpha=\beta=b=0,\quad\gamma=-\delta=1+\cos(\theta),\quad\mu=1.$
Then, for any $n\in\mathbb{N}$ and any $\lambda\in A_{n}(\theta)+B_{n}(\theta)$, there exists $\phi\in[0,2\pi)$ such that
$\lambda=\left\{\begin{array}[]{ll}2\sin\theta\cos\phi&\quad\sin\phi\neq 0,\\ 2\sin\theta&\quad\sin\phi=0,\end{array}\right$.
and $\phi$ satisfies
$\sin^{2}\theta\sin((2n+1)\phi)-(1+\cos\theta)^{2}\sin(2n-1)\phi=0.$
From Fact 2.5, the eigenvalues of $A_{n}(\theta)+B_{n}(\theta)$ are distinct for any $n\in\mathbb{N}$. Moreover, from statements (iv) and (v), statements (vi) and (vii) hold.
Next, we will prove (viii). Let $\lambda_{0}\in(-2\sin\theta,2\sin\theta)$ and $\eta>0$ such that $|\lambda_{0}|+\eta<2\sin\theta$. Define $r$, $J$, $2\pi]\to\mathbb{R}, R0R_{0},$ $C$ as follows:
$r$ $=\dfrac{\sin\theta}{1+\cos\theta},\quad R(\phi)=2((r-\cos\phi)^{2}+\sin^{2}\phi)\quad(\phi\in[0,2\pi]),$
$J$ $=\left\{\phi\in[0,2\pi]\middle|2\sin\theta\cos\phi\in[\lambda_{0}-\eta,\lambda_{0}+\eta]\right\},R_{0}=\inf\left(R\left(J\right)\right),$
$C$ $=\inf\left\{|1-\exp(i2\phi)|\ \middle|\phi\in J\right\}.$
Clearly $\sin\phi\neq 0$ for any $\phi\in J$, and $R_{0}$ and $C$ are positive.
Let $\lambda\in(\lambda_{0}-\eta,\lambda_{0}+\eta)$. In the present setting, for any eigenvalue $\lambda$ of $A_{n}(\theta)+B_{n}(\theta)$ and any eigenvector $u$ of $X$ corresponding to $\lambda$, there exists $\phi\in J$ such that
$u_{j}$ $=\dfrac{2u_{1}i}{\sqrt{\omega}}\left\{\sin\theta\sin(j\phi)-(1+\cos\theta)\sin((j-1)\phi)\right\}$
$=-\dfrac{2u_{1}i}{\sqrt{\omega}}(1+\cos\theta)(\sin((j-1)\phi)-r\sin(j\phi))\quad(j=1,2,\dots,2n).$ (5.1)
Then we have
$\quad-4\sum_{k=0}^{2n-1}(\sin(k\phi)-r\sin((k+1)\phi))^{2}$
$=\sum_{k=0}^{2n-1}(\exp(ik\phi)-\exp(-ik\phi)-r(\exp(i(k+1)\phi)-\exp(-i(k+1)\phi)))^{2}$
$=\sum_{k=0}^{2n-1}(\exp(ik\phi)-\exp(-ik\phi)-r\exp(i(k+1)\phi)+r\exp(-i(k+1)\phi))^{2}$
$=\sum_{k=0}^{2n-1}(\exp(i2k\phi)+\exp(i(-2k\phi))+r^{2}\exp(i2(k+1)\phi)+r^{2}\exp(-i2(k+1)\phi)$
$\quad-2-2r\exp(i(2k+1)\phi)+2r\exp(-i\phi)+2r\exp(i\phi)-2r\exp(-i(2k+1)\phi)$
$\quad-2r^{2})$
$=\sum_{k=0}^{2n-1}(\exp(i(2k\phi))+\exp(i(-2k\phi))+r^{2}\exp(i2(k+1)\phi)+r^{2}\exp(-i2(k+1)\phi)$
$\quad-2r\exp(i(2k+1)\phi)-2r\exp(-i(2k+1)\phi)-2(r^{2}-2(\cos\phi)r+1))$
$=-2nR(\phi)+\dfrac{\exp(i4n\phi)-1}{\exp(i2\phi)-1}+\dfrac{\exp(-i4n\phi)-1}{\exp(-i2\phi)-1}+r^{2}\exp(i\phi)\dfrac{\exp(i4n\phi)-1}{\exp(i2\phi)-1}$
$\quad+r^{2}\exp(-i\phi)\dfrac{\exp(-i4n\phi)-1}{\exp(-i2\phi)-1}-2r\exp(i\phi)\dfrac{\exp(i4n\phi)-1}{\exp(i2\phi)-1}$
$\quad-2r\exp(-i\phi)\dfrac{\exp(-i4n\phi)-1}{\exp(-i2\phi)-1}$
$=-2nR(\phi)+2\operatorname{Re}\left((\exp(i\phi)r^{2}-2\exp(i\phi)r+1)\dfrac{\exp(i4n\phi)-1}{\exp(i2\phi)-1}\right).$
Hence, for any sufficiently large $n\in\mathbb{N}$
$\left\|-2nR(\phi)+2\operatorname{Re}\left((\exp(i\phi)r^{2}-2\exp(i\phi)r+1)\dfrac{\exp(i4n\phi)-1}{\exp(i2\phi)-1}\right)\right\|$
$\geq 2nR_{0}-\dfrac{4(r+1)^{2}}{C}.$
It follows that $\dfrac{|u_{2n}|}{\|u\|}\to 0\quad\left(n\to\infty\right)$. Consequently, statement (viii) holds. ∎
The spectral radius of $A(\theta),B(\theta)]$ for $\theta\in(0,\pi)$ is given below.
Example 5.5.
The following statements hold:
  1. (i)
    The spectrum of $A(\theta)+B(\theta)$ is given by
    $(-2\sin\theta,2\sin\theta)\subset F(A(\theta)+B(\theta))\subset\sigma(A(\theta)+B(\theta))=[-2\sin\theta,2\sin\theta].$
  2. (ii)
    The following equality holds:
    $\sup_{\lambda\in F(A(\theta)+B(\theta))}b(\lambda)=\lim_{n\to\infty}b(\lambda_{n})=\rho([A(\theta),B(\theta)]),$
    where $\lambda_{n}$ is the eigenvalue of $A_{n}(\theta)+B_{n}(\theta)$ whose absolute value is closest to $\sqrt{2}$ for $n\in\mathbb{N}$.
  3. (iii)
    For any $\theta_{1}\neq\theta_{2}\in(0,\pi)$, the unitary representation $\tau_{P(\theta_{1}),Q(\theta_{1})}$ is not unitarily equivalent to the unitary representation $\tau_{P(\theta_{2}),Q(\theta_{2})}$.
  4. (iv)
    We have
    $\rho([A(\theta),B(\theta)])=\left\{\begin{array}[]{ll}2|\sin(2\theta)|&\quad|\theta-\dfrac{\pi}{2}|>\dfrac{\pi}{4},\vskip 5.69054pt\\ 2&\quad|\theta-\dfrac{\pi}{2}|\leq\dfrac{\pi}{4}.\end{array}\right$.
Proof of (i).
Let $\lambda\in(-2\sin\theta,2\sin\theta)$. From statements (vi) and (vii) of Lemma 5.4, there exists $\{\eta_{n}\}_{n=1}^{\infty}\subset(-2\sin\theta,2\sin\theta)$ and $\{u_{n}\}_{n=1}^{\infty}$ such that $\lim_{n\to\infty}\eta_{n}=\lambda$, and for any $n\in\mathbb{N}$, $u_{n}$ is a unit eigenvector of $A_{n}(\theta)+B_{n}(\theta)$ corresponding to $\eta_{n}$. Then for any $n\in\mathbb{N}$, $P_{n}(\theta)u_{n}=P(\theta)u_{n},\|(Q_{n}(\theta)-Q(\theta))u_{n}\|\leq 2|u_{n,2n}|$, where $u_{n,2n}$ is the $2n$-th component of $u_{n}$. From statement (viii) of Lemma 5.4, we have $\lim_{n\to\infty}\|(Q_{n}(\theta)-Q(\theta))u_{n}\|=0$. It follows that $\lambda\in F(A(\theta)+B(\theta))$. Hence we have $(-2\sin\theta,2\sin\theta)\subset F(A(\theta)+B(\theta))$. Since $\sigma(A(\theta)+B(\theta))$ is closed and $F(A(\theta)+B(\theta))\subset\sigma(A(\theta)+B(\theta))$, we conclude $2\sin\theta,2\sin\theta]\subset\sigma(A(\theta)+B(\theta))$.
We will prove that $\sigma(A(\theta)+B(\theta))\subset[-2\sin\theta,2\sin\theta]$. Let $\lambda\in\sigma(A(\theta)+B(\theta))$. Since $A(\theta)+B(\theta)$ is self-adjoint, $\lambda$ is in the approximate point spectrum of $A(\theta)+B(\theta)$. Therefore, there exists a sequence $\{u_{i}\}_{i=1}^{\infty}$ of unit vectors in $V$ such that $\|(A(\theta)+B(\theta))u_{i}-\lambda u_{i}\|<\dfrac{1}{2i}$ for any $i\in\mathbb{N}$. Since $\{(V_{i},\{P_{i}(\theta),Q_{i}(\theta)\})\}_{i=1}^{\infty}$ is a finite dimensional approximation of $(\ell^{2},\{P(\theta),Q(\theta)\})$, there exists a sequence $\{v_{i}\}_{i=1}^{\infty}$ of unit vectors in $V$ and exists a sequence $\varphi(1)<\ldots<\varphi(i)<\ldots$ of natural numbers such that $v_{i}\in V_{\varphi(i)}$ and $\|(A(\theta)+B(\theta))v_{i}-\lambda v_{i}\|<\dfrac{1}{i}$ for any $i\in\mathbb{N}$. It follows that $((A(\theta)+B(\theta))v_{i},v_{i})-\dfrac{1}{i}\leq\lambda\leq((A(\theta)+B(\theta))v_{i},v_{i})+\dfrac{1}{i}$ for any $i\in\mathbb{N}$. From Lemma 5.4, for any $i\in\mathbb{N}$
$((A(\theta)+B(\theta))v_{i},v_{i})=\lim_{n\to\infty}((A_{n}(\theta)+B_{n}(\theta))v_{i},v_{i})\in[-2\sin\theta,2\sin\theta].$
Thus, $2\sin\theta-\dfrac{1}{i}\leq\lambda\leq 2\sin\theta+\dfrac{1}{i}$ for any $i\in\mathbb{N}$. Letting $i\to\infty$, we have $2\sin\theta\leq\lambda\leq 2\sin\theta$. ∎
Proof of (ii).
From statements (vi)-(viii) of Lemma 5.4, we have
$\limsup_{n\to\infty}b(\lambda_{n})\leq\sup_{\lambda\in F(A(\theta)+B(\theta))}b(\lambda).$
Moreover, from statement (ii) of Theorem 5.3, we have
$\sup_{\lambda\in F(A(\theta)+B(\theta))}b(\lambda)=\lim_{n\to\infty}b(\lambda_{n})=\rho([A(\theta),B(\theta)]).$
Proof of (iii).
This follows directly from statement (i) and the fact that for any unitary representation $(\tau,V)$ of $D_{\infty}$ the spectrum of $\tau(g_{1})+\tau(g_{2})$ is invariant under any unitary intertwining operator. ∎
Proof of (iv).
From statement (i), we have
$\sup_{\lambda\in F(A(\theta)+B(\theta))}b(\lambda)=\left\{\begin{array}[]{ll}2|\sin(2\theta)|&\quad|\theta-\dfrac{\pi}{2}|>\dfrac{\pi}{4},\vskip 5.69054pt\\ 2&\quad|\theta-\dfrac{\pi}{2}|\leq\dfrac{\pi}{4}.\end{array}\right$.
Moreover, from statement (ii), statement (iv) holds. ∎
5.3 Application: Estimating spectral radius of Bell-CHSH operator
For bipartite quantum systems, we use the following notation.
Notation 5.6.
Let
  1. (S1)
    $V_{1}$ and $V_{2}$ be separable Hilbert spaces,
  2. (S2)
    $P_{i,j}$ be orthogonal projections on $V_{i}$ for $i,j\in\{1,2\}$,
  3. (A1)
    $(V_{i},\{P_{i,1},P_{i,2}\})$ satisfies either Case 1 or Case 2 for any $i\in\{1,2\}$, and
  4. (S3)
    for $i\in\{1,2\}$, let $\mathcal{A}_{i}=\{(V_{i,n},\{P_{i,1,n},P_{i,2,n}\})\}_{n=1}^{\infty}$ be a finite dimensional approximation of $(V_{i},\{P_{i,1},P_{i,2}\})$,
  5. (S4)
    for $i\in\{1,2\}$, when $(V_{i},\{P_{i,1},P_{i,2}\})$ satisfies Case 1, let $\mathcal{A}_{i}$ be the approximation defined in Lemma 5.2,
  6. (S5)
    for $i\in\{1,2\}$, when $(V_{i},\{P_{i,1},P_{i,2}\})$ satisfies Case 2, let $\mathcal{A}_{i}$ be the approximation defined in Theorem 4.6.
Under these assumptions, we define $A_{i},B_{i},A_{i,n},B_{i,n},\lambda_{i,n}$ $(i\in\{1,2\},n\in\mathbb{N})$ as follows:
  1. (i)
    $A_{i}=2P_{1,i}-\operatorname{id}_{V_{1}},B_{i}=2P_{2,i}-\operatorname{id}_{V_{2}}$,
  2. (ii)
    $A_{i,n}=2P_{1,i,n}-\operatorname{id}_{V_{1,n}}$,
  3. (iii)
    $B_{i,n}=2P_{2,i,n}-\operatorname{id}_{V_{2,n}}$,
  4. (iv)
    $\lambda_{1,n}$ is the eigenvalue of $A_{1,n}+A_{2,n}$ whose absolute value is closest to $\sqrt{2}$,
  5. (v)
    $\lambda_{2,n}$ is the eigenvalue of $B_{1,n}+B_{2,n}$ whose absolute value is closest to $\sqrt{2}$.
From Theorem 5.3, we obtain upper and lower bounds for the spectral radius of the Bell-CHSH operator, as follows.
Theorem 5.7.
Assume the setups (S1)-(S5), assumption (A1), notation (i)-(v) in Notation 5.6. Under these assumptions, we have
$\sup_{\begin{subarray}{c}\lambda\in\tilde{F}(A_{1}+A_{2}),\\ \lambda^{\prime}\in\tilde{F}(B_{1}+B_{2})\end{subarray}}\sqrt{4+b(\lambda)b(\lambda^{\prime})}\leq\rho(\mathcal{B})\leq\liminf_{n\to\infty}\sqrt{4+b(\lambda_{1,n})b(\lambda_{2,n})}.$
Acknowledgements
Y. F. would like to acknowledge Koichi Tojo, Takayuki Okuda, Yoji Koyama, Kohei Hatano, and Yuichiro Tanaka for their valuable comments and suggestions.